### v1

# Light-Weight Diffusion Multiplier and Uncertainty Quantification for Fourier Neural Operators

Abstract
Operator learning is a powerful paradigm for solving partial differential equations, with Fourier Neural Operators serving as a widely adopted foundation. However, FNOs face significant scalability challenges due to overparameterization and offer no native uncertainty quantification – a key requirement for reliable scientific and engineering applications. Instead, neural operators rely on post hoc UQ methods that ignore geometric inductive biases. In this work, we introduce DINOZAUR: a diffusion-based neural operator parametrization with uncertainty quantification. Inspired by the structure of the heat kernel, DINOZAUR replaces the dense tensor multiplier in FNOs with a dimensionality-independent diffusion multiplier that has a single learnable time parameter per channel, drastically reducing parameter count and memory footprint without compromising predictive performance. By defining priors over those time parameters, we cast DINOZAUR as a Bayesian neural operator to yield spatially correlated outputs and calibrated uncertainty estimates. Our method achieves competitive or superior performance across several PDE benchmarks while providing efficient uncertainty quantification.
1 Introduction
Partial differential equations (PDEs) are a fundamental mathematical tool that describe a wide range of physical phenomena observed in the real world. Numerous problems in science and engineering require formulating and solving PDEs, often presenting significant computational and methodological challenges to researchers and practitioners. A common strategy for solving PDEs is via conventional numerical methods, such as finite difference, finite element, or finite volume method Larsson and Thomée [2003]. While accurate, they become prohibitively expensive in complex engineering tasks, especially where rapid iteration or real-time response is critical, such as design optimization or control Martins and Ning [2021].
In recent years, there has been a growing trend toward replacing computationally intensive numerical solvers with neural network-based surrogate models trained to approximate PDE solutions. Since discretized PDE domains are commonly represented using point clouds, meshes, or graphs, geometric deep learning Bronstein et al. [2021] has emerged as a promising paradigm for this class of problems. In particular, message passing layers, initially developed for geometric and relational data, have proven effective due to their structural resemblance to physical convolution operators. These layers are especially suitable for capturing the spatial dependencies inherent in PDE systems, making them well-suited for learning dynamics over irregular or non-Euclidean domains Pfaff et al. [2020], Brandstetter et al. [2022].
However, geometric deep learning methods often lack key properties required for solving PDEs, motivating alternative approaches. A PDE – its parameters, equation structure, and boundary conditions – defines an operator that maps input functions to solution functions. Neural operators (NOs) Lu et al. [2019], Kovachki et al. [2023, 2024] have recently emerged as a leading framework for approximating such solution operators and, more generally, for tackling the operator learning problem implicitly present in PDEs. These models incorporate inductive biases that improve performance and offer discretization invariance and universal approximation Kovachki et al. [2023], making them a strong alternative to geometric deep learning. A prominent example is the Fourier Neural Operator (FNO) Li et al. [2021], which learns transformations in the spectral domain via the Fast Fourier Transform. While FNOs are restrictive in requiring input data on a regular grid, they remain a widely used backbone in neural operator architectures Li et al. [2024]. Despite their success on problems such as Darcy flow and Burgers’ equation, FNOs are known to be overparameterized Tran et al. [2023], Kossaifi et al. [2024]. Their design leads to exponential growth with domain dimensionality, causing overfitting and significant memory overhead.
Figure 1: Overview of DINOZAUR. NO block $\mathcal{N}_{\theta_{i},\tau_{i}}$ is revised by updating the integral transform $\mathcal{I}_{\tau_{i}}$ with diffusion multiplier $\exp(-\lambda_{k}\tau_{i})$ and including gradient features $\mathcal{G}_{\theta_{i}}$ to add anisotropy. Feed-forward network (FFN) is applied to mix two sets of features.
Moreover, PDE solvers often inform downstream decision-making, so it is imperative for any substitute method to quantify the uncertainty on its output. Uncertainty quantification (UQ) is a common strategy to improve model robustness by estimating confidence in its predictions. It enhances model outputs by accounting for different sources of error, helping guide more informed and reliable downstream use. Uncertainty modeling has long relied on Bayesian inference Mackay [1992], ensembles Dong et al. [2020], and conformal prediction Shafer and Vovk [2008] to produce confidence estimates. While some of these methods have been adapted to deep learning with varying success Gal and Ghahramani [2016], Zhang et al. [2021], Jospin et al. [2022], UQ for GNNs and neural operators Psaros et al. [2023] remains underdeveloped, often relying on post hoc methods ill-suited to their structure.
To address the scalability issues and limited UQ methods in FNOs, we present DINOZAUR – a diffusion-based neural operator parametrization with uncertainty quantification. Our diffusion multiplier is inspired by the heat kernel in the Fourier basis and requires only one learnable parameter per channel, resulting in superior scalability without sacrificing predictive performance. This physically motivated choice of multiplier allows us to place meaningful probabilistic priors on the learnable time parameters, recover the posterior via inexpensive variational inference (VI) Blei et al. [2017] of these parameters, and efficiently sample spatially correlated functions. We summarize our contributions as follows:
  • We introduce dimensionality-independent spectral multiplier parametrization, reducing the parameter count by several orders of magnitude compared to equivalent FNO-based architectures.
  • We define a Bayesian formulation of the heat kernel and build an interpretable uncertainty quantification method to sample spatially correlated outputs.
  • We demonstrate that our method is scalable, maintains strong predictive performance across multiple PDE benchmarks, and achieves superior UQ estimates compared to conventional probabilistic methods.
2 Related work
The area of designing learnable approximation for PDE solutions is vast and rapidly growing Wang et al. [2024], Huang et al. [2025]. Given the increasing availability of data, the need to incorporate inductive biases, and the inherently geometric structure of many physical problems, two classes of deep learning models have come to dominate the field: geometric deep learning models and neural operators.
Geometric deep learning
Graph neural networks (GNNs) Scarselli et al. [2008] offer a flexible framework for learning on non-Euclidean domains. Many GNNs implement spatial convolution by performing learnable aggregation of local neighborhoods Niepert et al. [2016], Chamberlain et al. [2021]. Neural PDE solvers increasingly rely on spatial GNN layers to model and predict physical phenomena across various domains, including control Li et al. [2019], deformable material simulation Sanchez-Gonzalez et al. [2020], and spatiotemporal mesh evolution Pfaff et al. [2020]. Specific architectures target building robust autoregressive solvers Brandstetter et al. [2022] and encoding the geometric boundary conditions Mayr et al. [2023]. However, their performance often suffers from sensitivity to discretization, limiting generalization across mesh resolutions. Spectral GNNs define convolutions using the graph Laplacian’s eigen-decomposition Bruna et al. [2014], Kipf and Welling [2017], often with polynomial approximations of spectral filters Defferrard et al. [2016], Levie et al. [2018]. These have seen success in 3D shape analysis, particularly with functional maps Ovsjanikov et al. [2012]. DiffusionNet Sharp et al. [2022] exemplifies this approach by using the spectral form of the heat kernel for a simple and interpretable kernel parametrization. It exhibits strong performance and generalization; we note that such a kernel is applicable in arbitrary domains.
Neural operators
Neural operators are a powerful class of models for learning mappings between infinite-dimensional spaces in a resolution-invariant manner Kovachki et al. [2023], offering an alternative to GNNs. Lu et al. [2019] introduced DeepONet, inspired by operator learning and universal approximation theory, using a branch-trunk architecture to encode inputs and evaluate the solution at arbitrary locations. A parallel line of work introduced the Graph Neural Operator (GNO) architecture Anandkumar et al. [2020], Li et al. [2020], transferring the message passing intuition from GNNs to NOs. A family of neural operators decomposes the input function into a basis defined by one of the popular transforms: Fourier Li et al. [2021], Laplace Cao et al. [2024], Wavelet Tripura and Chakraborty [2023]. In particular, FNOs are popular for their performance and interpretability despite needing a regular grid sampling, and are commonly used in other networks, such as Geo-FNO Li et al. [2023], U-FNO Wen et al. [2022], and Geometry-Informed Neural Operator (GINO) Li et al. [2024], to name a few. However, FNOs face scalability and generalization issues due to exponential parameter growth with domain dimensionality Tran et al. [2023]. Efforts to improve this, such as FFT factorization Tran et al. [2023], low-rank tensor decomposition Kossaifi et al. [2024], and MLPs in the spectral domain Xiao et al. [2024], have reduced growth but still rely on spatial dimensions.
Functional uncertainty quantification
UQ in deep learning has been widely studied, with popular methods including MC Dropout Gal and Ghahramani [2016], Laplace approximation MacKay [1992], ensembles Rahaman et al. [2021], and non-parametric conformal prediction Zhang et al. [2021], which typically require extra computations or careful calibration. In the context of functional UQ – crucial for neural operators – the landscape is similarly limited. Recent approaches include applying a post hoc Laplace approximation to the final layer of the NO Magnani et al. [2022], using conformal methods after training Ma et al. [2024], or learning an uncertainty measure as an auxiliary output Bülte et al. [2025]. However, none of these strategies model uncertainty within the NO; instead, they adopt generic methods for neural networks and neglect the inductive biases of neural operators.
3 Diffusion multiplier parametrization for neural operators
3.1 Problem setting
We consider the task of learning a parametric surrogate to map the initial and boundary conditions of a PDE to its solution. Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set representing a PDE domain in $d$-dimensional Euclidean space. Let also $\mathcal{A}=\mathcal{A}(\Omega;\mathbb{R}^{d_{a}})$ be a Banach space of real-valued functions defined on $\Omega$ and $\mathcal{U}=\mathcal{U}(\Omega;\mathbb{R}^{d_{u}})$ be a space of solution functions. We associate $a\in\mathcal{A}$ with a function that parametrizes the partial differential operator $\mathcal{L}_{a}$ and trace operator $\mathcal{B}_{a}$. Then, we form the PDE:
$\mathcal{L}_{a}[u](x)$ $=f(x)\quad\forall x\in\Omega,$ (1)
$\mathcal{B}_{a}[u](x)$ $=g(x)\quad\forall x\in\partial\Omega\,,$
where $f(x)$ is a fixed forcing function and $g(x)$ is a boundary condition. This formulation is general enough to include varying geometric boundaries by including the signed distance function of a geometry in functions from $\mathcal{A}$ and defining a solution as an extension to the whole domain Stein [1970], Li et al. [2024].
We assume the PDE induces a continuous solution map $\mathcal{N}:\mathcal{A}\rightarrow\mathcal{U}$ that maps inputs $a$ to solutions $u$, and we are seeking to find its parametrized approximation $\mathcal{N}_{\theta}$ with $p$ parameters. Given the discretized observations, $\{a_{n},u_{n}\}_{n=1}^{N}$, we pose a problem of the empirical risk minimization :
$\min_{\theta\in\mathbb{R}^{p}}\frac{1}{N}\sum_{n=1}^{N}\|u_{n}-\mathcal{N}_{\theta}[a_{n}]\|_{\mathcal{U}}^{2}.$ (2)
We construct the neural operator in the usual form as a sequence of lifting layer $\mathcal{P}$, $M$ NO blocks $\mathcal{N}_{\theta_{i}}$, and projection layer $\mathcal{Q}$, where the NO block $\mathcal{N}_{\theta_{i}}$ is given by:
$v_{i+1}(x)=\mathcal{N}_{\theta_{i}}[v_{i}](x)=\sigma\left(W^{\mathrm{skip}}_{i}v_{i}(x)+b_{i}+\int_{\Omega}\kappa_{\gamma_{i}}(x,y)v_{i}(y)\mathrm{d}y\right),\,i=1,\dots,M,$ (3)
where $\kappa_{\gamma_{i}}$ is a kernel, matrix $W^{\mathrm{skip}}\in\mathbb{R}^{d_{c}\times d_{c}}$, $b_{i}\in\mathbf{R}^{d_{c}}$ is a bias and $d_{c}$ is the number of channels. FNO restricts domain to be periodic torus $\Omega=\mathbb{T}^{d}$, kernels to be stationary and represents the integral transform $\mathcal{I}$ in the following way:
$\mathcal{I}_{\gamma}[v](x)=\int_{\Omega}\kappa_{\gamma}(x-y)v(y)\mathrm{d}y=\mathcal{F}^{-1}\bigl{[}R_{\gamma}(k)\mathcal{F}[v](k)\bigr{]}(x),$ (4)
where $\mathcal{F}$ is Fourier transform, and $R_{\gamma}(k)$ acts as a unique complex-valued matrix if $k<k_{\mathrm{max}}\in\mathbb{Z}^{d}$ and as zero otherwise.
Given the block’s width $d_{c}$ and maximal truncation modes $k_{\mathrm{max}}=(k^{1}_{\mathrm{max}},\dots,k^{d}_{\mathrm{max}})\in\mathbb{Z}^{d}$, tensor $R_{\gamma}=[R_{\gamma}(k)]_{k\leq k_{\mathrm{max}}}$ contains $O(d_{c}^{2}\cdot\Pi_{j=1}^{d}k_{\mathrm{max}}^{j})$ learnable parameters, yielding quadratic dependency on the number of channels and exponential growth of parameters in the dimensionality $d$ of the domain $\Omega$. In this paper, we target the form of kernel $\kappa_{\gamma}$ and propose an alternative parametrization.
3.2 DINOZAUR architecture
Diffusion multiplier
Motivated by the success of DiffusionNet Sharp et al. [2022] and its resemblance to FNO, we choose to restrict the form of $\kappa_{\gamma}$ to that of the heat kernel for our architecture. Consider the heat equation on the torus $\Omega=\mathbb{T}^{d}$:
$\partial_{\tau}v(\tau,x)$ $=\Delta v(\tau,x)$ (5)
$v(0,x)$ $=v_{0}(x),$
with a solution Strauss [2007] given by:
$v(\tau,x)=\mathcal{I}_{\tau}[v_{0}](x)=\mathcal{F}^{-1}\bigl{[}R_{\tau}(k)\mathcal{F}[v_{0}](k)\bigr{]}(x)=\mathcal{F}^{-1}\bigl{[}\exp(-\lambda_{k}\tau)\odot\mathcal{F}[v_{0}](k)\bigr{]}(x),$ (6)
where $\odot$ is the Hadamard product and $\lambda_{k}=-4\pi^{2}\|k\|^{2}$. We refer to $R_{\tau}=[\exp(-\lambda_{k}\tau)]_{k\leq k_{\mathrm{max}}}$ as a diffusion multiplier. Figure 1 illustrates the updated block architecture.
When integrated into the FNO block, $R_{\tau}\in\mathbb{R}^{k^{1}_{\mathrm{max}}\times\dots\times k^{d}_{\mathrm{max}}\times d_{c}}$ is a tensor with the only learned parameters being $\tau=(\tau^{1},\dots,\tau^{d_{c}})\in\mathbb{R}^{d_{c}}$. Each $\tau^{c}$ is an independent scalar defining diffusion time within channel $c$ of the NO block. Times $\tau$ control the adjustable receptive field of the convolution kernel and are not related to the time variable that may be present in the original PDE (1). We note that our multiplier parametrization does not depend on dimensionality $d$ of the domain $\Omega$ or on the truncation modes $k_{\mathrm{max}}$ since $\lambda_{k}$ is precomputed and fixed.
Gradient features
The diffusion multiplier propagates information across the domain, enabling efficient function convolution. However, the heat kernel is isotropic by construction, which may limit its applicability. We found that including gradient features mitigates that issue and yields better experimental results.
Since we already map to Fourier basis, it’s natural to calculate gradients in the spectral domain. Raw gradients can be ill-defined due to non-periodic boundaries or uneven sampling. Following Sharp et al. [2022], to improve robustness and add some flexibility, we apply a linear transformation $W^{\mathrm{grad}}\in\mathbb{R}^{d_{c}\times d_{c}}$ with a hyperbolic tangent non-linearity to normalize the output in the physical domain. We denote the gradient features operator as $\mathcal{G}_{\theta}$ and define its output at channel $c$ as follows:
$\nabla v(x)$ $=\mathcal{F}^{-1}\bigl{[}i2\pi k\mathcal{F}[v](k)\bigr{]}(x)$ (7)
$\mathcal{G}_{\theta}[v]^{c}(x)$ $=\operatorname{tanh}\left(\sum_{r=1}^{d_{c}}\Bigl{\langle}\nabla v^{c}(x),W^{\mathrm{grad}}_{cr}\nabla v^{r}(x)\Bigr{\rangle}\right),\,c=1,\dots,d_{c}.$
Since $v\in\mathbb{R}^{d_{c}}$, its Jacobi matrix is $\nabla v\in\mathbb{R}^{d_{c}\times d}$. By taking the inner product, we reduce the spatial dimensionality $d$ and end up with a compatible shape. We report the results with and without the gradient features in Section 4.1.
Gradient features are applied to the output of the integral transform $\mathcal{I}_{\tau}$ and stacked with it inside the DINOZAUR block, doubling the number of channels. We add a matrix $W^{\mathrm{mix}}\in\mathbb{R}^{d_{c}\times 2d_{c}}$ to map all features back to the width of the block. The final block architecture is:
$v_{i+1}(x)=\sigma\left(W^{\mathrm{skip}}_{i}v_{i}(x)+b_{i}+W_{i}^{\mathrm{mix}}\begin{bmatrix}\mathcal{I}_{\tau_{i}}[v_{i}](x)\\ \big{(}\mathcal{G}_{\theta_{i}}\circ\mathcal{I}_{\tau_{i}}\big{)}[v_{i}](x)\end{bmatrix}\right).$ (8)
Substituting the original blocks in (3) in the neural operator $\mathcal{N}_{\theta}$ with modified blocks from (8) results in DINOZAUR architecture that we denote as $\mathcal{N}_{\theta,\tau}$, with all non-diffusion parameters lumped together in $\theta$ and separated from $\tau$.
Universal approximation
Despite the seeming simplicity of the diffusion multiplier, we claim that our formulation of a neural operator exhibits universal approximation properties: see Proposition 1.
Proposition 1.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary and such that the closure $\overline{\Omega}\subset(0,2\pi)^{d}$. For given integers $s,s^{\prime}$ > 0, let $\mathcal{N}:C^{s}(\overline{\Omega};\mathbb{R}^{d_{a}})\rightarrow C^{s^{\prime}}(\overline{\Omega};\mathbb{R}^{d_{u}})$ be a continuous operator, and fix a compact set $\mathbf{A}\subset C^{s}(\overline{\Omega};\mathbb{R}^{d_{a}})$. Then there exists a continuous, linear operator $\mathcal{E}:C^{s}(\Omega;\mathbb{R}^{d_{a}})\rightarrow C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})$ such that $\mathcal{E}[a]|_{\Omega}=a$ for all $a\in C^{s}(\Omega;\mathbb{R}^{d_{a}})$. Furthermore, for any $\varepsilon>0$, there exists a DINOZAUR such that
$\sup_{a\in\mathbf{A}}\|\mathcal{N}[a]-(\mathcal{N}_{\theta,\tau}\circ\mathcal{E}[a])|_{\Omega}\|_{C^{s^{\prime}}(\overline{\Omega};\mathbb{R}^{d_{u}})}\leq\varepsilon,$
where $\mathcal{N}_{\theta,\tau}=\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\tau}\circ\mathcal{P}:C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})\rightarrow C^{s^{\prime}}(\mathbb{T}^{d},\mathbb{R}^{d_{u}})$.
We provide the proof in the Appendix E.
3.3 Bayesian inference for diffusion multiplier
$a$$u$$\tau$$\sigma$$\theta$ $N$
Figure 2: (a) Probabilistic graph for DINOZAUR, showing how the latent diffusion-times, $\tau$, relate to the observations and non-mechanistic model parameters, $\theta,\sigma$. (b) We define priors for $\tau$ and sample multiple time parameters from posterior in each block at inference time, which translates into spatially correlated uncertainty at the output.
Now that we established the overall architecture, we will describe the Bayesian inference scheme to obtain UQ. An advantage of the proposed diffusion multiplier as a message passing scheme is the ability to impose meaningful priors on the diffusion-time parameters $\tau$. Doing so recasts DINOZAUR as a Bayesian neural operator, with the source of uncertainty coming from the integral transform. In our method, other model parameters $(\theta,\sigma)$ are treated deterministically. The conciseness of our multiplier’s parametrization allows for tractable inference and alleviates the need to grapple with extremely high-dimensional distributions. We define our conditional generative model for any output signal $u_{n}$ given an input signal $a_{n}$ as
$\ln\tau^{c}_{i}$ $\sim\mathbf{N}(\mu_{\text{prior}},\sigma^{2}_{\text{prior}}),\,i=1,\dots,M;\,c=1,\dots,d_{c}$ (9)
$u_{n}(\cdot)|a_{n}(\cdot),\tau$ $\sim\mathbf{N}(\mathcal{N}_{\theta,\tau}[a_{n}](\cdot),\sigma^{2}),$
where $f(\cdot)$ denotes functions evaluated at any finite number of points in the domain $\Omega$, $\mathbf{N}$ denotes a normal distribution, and $\mu_{\text{prior}},\sigma_{\text{prior}}$ are hyperparameters of the prior distribution. We collect all time variables from each channel $c$ and block $i$ for notational convenience, $\tau=\left\{(\tau^{c}_{i})_{c=1}^{d_{c}}\right\}_{i=1}^{M}$. See Figure 2(a) for the corresponding graphical model. We also collect training data in a set of discrete observations, $\mathcal{D}=\left\{\{a_{nk}\}_{k=1}^{K_{n}},\{u_{nj}\}_{j=1}^{J_{n}}\right\}_{n=1}^{N}$, with possibly different sets of evaluation points $\{x_{k}\}$, $\{y_{j}\}$.
We seek to recover the approximate posterior for the diffusion time, $p_{\theta,\sigma}(\tau|\mathcal{D})$, according to (9), which yields a posterior for the diffusion multipliers, $R_{\tau}$, and the entire neural operator $\mathcal{N}_{\theta,\tau}$. This can be used to produce a posterior predictive distribution over the output for a new input function $a_{\star}$ (see Figure 2(b)):
$p_{\theta,\sigma}(u_{\star}(\cdot)|a_{\star},\mathcal{D})=\int p_{\theta,\sigma}(\tau|\mathcal{D})\mathbf{N}(u_{\star}(\cdot);\mathcal{N}_{\theta,\tau}[a_{\star}](\cdot),\sigma^{2})\mathrm{d}\tau,$ (10)
where we marginalize over $\tau$ to reflect epistemic uncertainty over the possible operators that fit the data, as well as aleatoric uncertainty from the assumed normal noise with fitted $\sigma$. Note that while the noise is point-wise independent for a fixed $\tau$, the $\tau$ dependence spatially correlates predictions.
Variational inference
As mentioned, the posterior $p_{\theta,\sigma}(\tau|\mathcal{D})$ is intractable, so we resort to VI to recover an approximate posterior. We assume that it can be adequately approximated by a distribution $q(\tau;\phi)$ parametrized by $\phi$. We choose $q(\tau;\phi)$ such that it factorizes over the $i=1,\dots,M$ blocks of the NO but keeps correlations between diffusion times of channels in a block:
$q(\tau;\phi)\dot{=}\prod_{i=1}^{M}q(\tau_{i};\phi_{i})\dot{=}\prod_{i=1}^{M}\mathbf{N}(\ln\tau_{i};\mu_{i},\Sigma_{i}),$ (11)
where $\tau_{i},\mu_{i}\in\mathbb{R}^{d_{c}}$, and $\Sigma_{i}\in\mathbb{R}^{d_{c}\times d_{c}}$. We collect the parameters of the approximating factors, $\phi=\{\phi_{i}\}_{i=1}^{M}=\{(\mu_{i},\Sigma_{i})\}_{i=1}^{M}$, for notational convenience. Due to their parsimonious and mechanistic role in controlling message propagation, diffusion time parameters avoid the redundancy and symmetries that challenge VI for typical neural network weights, allowing for better identifiability and justifying a full-rank approximation within each block.
We then seek to find $\phi$ that minimizes the Kullback-Leibler divergence (KL) from the $q(\tau;\phi)$ to the true posterior. As this objective is still intractable Murphy [2022], we instead maximize a lower bound (ELBO):
$\mathcal{L}_{\rm ELBO}(\theta,\sigma;\phi)=$ $\mathbb{E}_{q(\tau;\phi)}[\ln p_{\theta,\sigma}(\mathcal{D}|\tau)]-D_{\rm KL}[q(\tau;\phi)\|p(\tau)],$ (12)
which allows to learn jointly the network parameters $(\theta,\sigma)$ and variational parameters $\phi$. During inference, we replace the true posterior $p_{\theta,\sigma}$ with its approximation $q(\tau;\phi)$ in (10).
4 Experiments
Table 1: Comparison of DINOZAUR against the baselines in deterministic setting, reporting the $\mathrm{RL}_{2}$ values on test split. The best result is bold, second best is $\underline{\text{underlined}}$
Model # params $\mathrm{RL}_{2}\downarrow$ params $\mathrm{RL}_{2}\downarrow$
(×$(\times 10^{3})$ (×$(\times 10^{-2})$ (×$(\times 10^{3})$ (×$(\times 10^{-2})$
Darcy flow Navier-Stokes
FNO 361.4 $\underline{0.64}$ 12,650.8 $\underline{15.96}$
FNOg 371.9 0.86 12,692.2 17.24
TFNO 17.7 4.44 69.1 21.39
TFNOg 28.2 2.28 110.6 19.40
[…]
DINOZAUR (Ours) 28.0 0.39 109.7 15.86
ShapeNet car Ahmed bodies
GINO 49,608.6 8.30 49,612.9 8.51
GINOg 49,650.1 7.03 49,645.3 $\underline{8.07}$
TGINO 149.8 7.42 154.0 8.13
TGINOg 191.3 8.09 195.5 8.86
[…]
DINOZAUR (Ours) 180.4 $\underline{7.08}$ 184.6 7.96
We evaluate the performance of the DINOZAUR architecture on several standard benchmarks. Section 4.1 presents results for the deterministic variant of DINOZAUR. In Section 4.2, we enable the Bayesian version and assess its UQ capabilities, comparing it to established UQ methods for neural operators. Ablation studies and computational performance comparisons are detailed in Section 4.3. All models are implemented in PyTorch Paszke et al. [2017] and trained on a single NVIDIA A100 GPU for 150 epochs with AdamW Loshchilov and Hutter [2019] and One-Cycle scheduler Smith and Topin [2019] or AdamW Schedule-Free Defazio et al. [2024] with learning rate (in most cases) set to $10^{-3}$. See the Appendix B for details.
Benchmarks
We evaluate our method on two regularly sampled datasets from Li et al. [2021] and two datasets with irregular domains used in Li et al. [2024]. For the 2D Darcy flow, we learn a mapping from a scalar permeability field of resolution $241^{2}$ to a scalar pressure field, applying standard scaling to both. The 2D Navier-Stokes (NS) dataset contains $64^{2}$ spatial points and 20 time steps of simulations with viscosity $\nu=10^{-5}$ on a uniform grid. We learn to map the first 10 time steps of vorticity to the last 10, without applying data scaling. For further dataset and PDE details, we refer to Takamoto et al. [2022]. The irregular-domain benchmarks include two datasets for predicting surface pressure on 3D meshes: ShapeNet car Umetani and Bickel [2018] and Ahmed bodies Ahmed et al. [1984]. Following the protocol of Li et al. [2024], we evaluate the signed distance function (SDF) of each mesh on a $64^{3}$ uniform grid and apply standard scaling to the outputs during training. We learn a map from SDF values to scalar pressure on mesh vertices, with inlet velocity added to input in the Ahmed dataset. A summary of the benchmark datasets is provided in the Appendix A.
Baselines
[…]
$\mathrm{RL}_{2}=\frac{1}{N_{\mathrm{test}}}\sum_{n=1}^{N_{\mathrm{test}}}\frac{\|u_{n}(x)-\widehat{u}_{n}(x)\|}{\|u_{n}(x)\|}.$
We emphasize that the only structural change our model makes is the multiplier in FNO blocks.
Figure 3: Uncertainty predictions. Left: mean of pressure samples on a test mesh with a geodesic path traced along the top. Right, top: ground truth and 100 sampled predictions along the same path. Right, bottom: car profile with a qualitative visualization of the standard deviation of pressure samples along the outline. High uncertainty is near sharp field changes, e.g., between Points 0 and 10.
4.1 Results: deterministic performance
Table 1 reports evaluation results. DINOZAUR significantly outperforms all baselines on Darcy and remains competitive on other benchmarks. The parameter counts in the table highlight a key trend: as problem dimensionality increases, FNO’s parameter count grows by two orders of magnitude, whereas DINOZAUR scales only with network width. Although TFNO shares DINOZAUR’s favorable scaling, it suffers notable performance drops on uniform datasets. Interestingly, reducing the parameter count in the non-uniform setting improves GINO’s performance, further underscoring the overparameterization issue. In contrast, DINOZAUR maintains strong results across domains, demonstrating the effectiveness of diffusion-based message passing as a parameter-efficient alternative.
Table 2: Evaluation of the uncertainty output quality of DINOZAUR and baseline models. The best result is bold, second best is $\underline{\text{underlined}}$
[…]
(×$(\times 10^{-2})$
ShapeNet car
GINOD $\underline{7.38}$ 18.58 0.36 16.64
GINOL 10.38 $\underline{3.15}$ $\underline{0.27}$ 15.30
TGINOD 9.18 54.13 0.41 23.17
TGINOL 9.28 3.07 $\underline{0.27}$ $\underline{14.17}$
DINOZAURD (Ours) 6.92 16.00 0.36 16.05
DINOZAURL (Ours) 8.82 $\underline{3.15}$ $\underline{0.27}$ 15.30
DINOZAURB (Ours) 7.46 4.35 0.20 12.73
[…]
4.2 Results: uncertainty quantification
In this section, we assess the UQ performance of the Bayesian version of our model, DINOZAURB, introduced in Section 3.3, on the more challenging non-uniform data. As baselines, we consider MC Dropout Gal and Ghahramani [2016] (NOD) and Laplace approximation Magnani et al. [2022] (NOL). MC Dropout is a widely used, model-agnostic UQ method, which we apply by enabling 10% dropout in the Fourier blocks’ MLPs to emulate stochasticity in the message passing. The Laplace approximation follows Magnani et al. [2022], modeling uncertainty via a fitted distribution over the parameters of the final MLP layer $\mathcal{Q}$. The discussion of hyperparameter choice, including $\mu_{\text{prior}},\sigma_{\text{prior}}$ required for setting $\ln\tau$ distributions is in the Appendix B.
UQ results in Table 2 are obtained by estimating the mean and standard deviation from 100 posterior predictive distribution samples per test data instance. Metrics are computed using the Uncertainty Toolbox Chung et al. [2021]; full metric definitions are provided in the Appendix D. DINOZAURB matches or surpasses all baselines, achieving substantially lower negative log-likelihood (NLL) than dropout models and comparable results to Laplace approximation. It outperforms all baselines in terms of miscalibration area (MA) and interval score (IS). An important consideration is that the Laplace approximation requires a secondary computationally expensive step to obtain the Hessian of the loss function with respect to the network parameters after training. In contrast, DINOZAURB natively supports UQ, thereby providing accurate uncertainty estimates out of the box. See results in Figure 3.
4.3 Ablation studies
To evaluate the scalability of DINOZAUR, we conduct experiments varying model width, depth, and the number of truncation modes using the ShapeNet car dataset. As shown in Figure 4(a), test error (measured by RL2 loss) decreases with increasing hyperparameters, saturating near the values used in our main experiments. Further improvements may be limited by the input data resolution of $64^{3}$. Figure 4(b) confirms our scaling claims: depth increases parameter count almost linearly, width affects it quadratically, while modes do not affect the model size.
[…]
Figure 4: Scalability and efficiency of DINOZAUR. (a) Test RL2 error on ShapeNet car as width, depth, and modes are varied. (b) Model size under the same settings. (c) Average peak GPU memory usage across five train-predict cycles on synthetic data. (d) Average time under the same settings.
5 Discussion
The recent emergence of neural operators has introduced a powerful alternative to traditional, computationally expensive numerical methods for solving PDEs in downstream applications. Architectures such as the Fourier Neural Operator and the Geometry-Informed Neural Operator have demonstrated notable success in learning solution operators across various physical systems. However, these models rely on highly parameterized transformations in spectral space, which can hinder scalability and interpretability. Furthermore, FNOs and other common architectures fail to provide native support for uncertainty quantification, a critical requirement for reliable deployment in scientific and engineering applications. As a result, practitioners are often forced to apply generic UQ techniques post hoc, which fail to exploit the spatio-temporal inductive biases intrinsic to the underlying physical systems.
In this work, we introduced DINOZAUR, an FNO-based model featuring a parsimonious and physically motivated parameterization of the message passing mechanism. This design stems from the analytic solution to the heat equation and is augmented by spatial gradient features. By presenting learned diffusion time parameters, DINOZAUR defines a new diffusion multiplier that is dimensionality-independent and offers more favorable scalability than the original FNO. Through extensive evaluation on benchmarks, we showed that DINOZAUR consistently matches or exceeds the predictive performance of dense FNO and GINO architectures while using orders of magnitude fewer parameters. Leveraging this efficient parameterization, we introduce meaningful priors over diffusion time parameters, obtaining a Bayesian neural operator that delivers competitive UQ performance against classical deep learning UQ methods such as MC Dropout and Laplace approximation. To the best of our knowledge, DINOZAUR is the first Bayesian NO to explicitly define distributions in the integral transform.
Limitations and future work
In our investigations, we found that the primary limitation of the diffusion mechanism is its radial symmetry. We observed that diffusion alone struggles to train well. Our solution is to include gradient features, though at the cost of increased computational time. Investigating other strategies, like introducing anisotropy intrinsically, using the modifications proposed in Liu-Schiaffini et al. [2024], or mixing gradients with other features before applying inverse FFT overhead, are promising directions for future work. We found that our Bayesian formulation is sensitive to hyperparameter settings of the prior distributions. Customizing priors across network layers to capture different physical properties could improve the robustness and interpretability of the model. More broadly, exploring other parsimonious kernel parameterizations may offer both interpretability and improved generalization, especially in settings with limited data or strong physical priors. Finally, since our parameterization is independent of spatial dimensionality, it opens up opportunities for pretraining on lower-dimensional or simpler physical systems, with the potential to accelerate learning in more complex, higher-dimensional domains through transfer learning Hao et al. [2024].
We believe our work is a step forward to a deeper understanding of neural operators, advancing them toward becoming reliable tools for real-world scientific and engineering applications.
Appendix A Benchmarks
Table 3: Summary of the benchmark datasets
Dataset $d$ Input type Output type $N_{\mathrm{train}},N_{\mathrm{test}}$
Darcy flow 2 Regular [241;241][241;241][ 241 ; 241 ] Regular [241;241][241;241][ 241 ; 241 ] 1024, 1024
NS 2D+T 2+1 Regular [64;64;10][64;64;10][ 64 ; 64 ; 10 ] Regular [64;64;10][64;64;10][ 64 ; 64 ; 10 ] 1000, 200
ShapeNet car 3 Regular [64;64;64][64;64;64][ 64 ; 64 ; 64 ] Irregular [3,586][3{,}586][ 3 , 586 ] 561, 100
[…]
Table 4: Data transformations during training
Dataset Inputs Input scaler Padded input Target Target scaler
Darcy flow Permeability Standard [271;271][271;271][ 271 ; 271 ] Pressure Standard
[…]
ShapeNet car SDF None [72;72;72][72;72;72][ 72 ; 72 ; 72 ] Pressure Standard
Ahmed bodies SDF, Inlet velocity None [72;72;72][72;72;72][ 72 ; 72 ; 72 ] Pressure Standard
To evaluate our model and compare it against the baselines, we selected two datasets with regularly sampled domains and two datasets with non-uniform outputs.
Darcy flow
This dataset contains steady-state 2D solutions of the Darcy flow equation on the unit box, modeling the flow of fluid through a porous medium Li et al. [2021]. It includes samples with spatial resolution of $\times 241$ 241 points. The network received a scalar permeability field as input and was tasked to predict scalar pressure values. To ensure non-periodicity, we applied padding during training, increasing the resolution to $\times 271$ 271. A standard scaler was applied to both input and output fields. Train and test splits had 1024 samples each.
Navier-Stokes
The Navier–Stokes dataset includes simulations of incompressible, viscous fluid flow with constant density and viscosity $\nu=10^{-5}$ Li et al. [2021]. It features a periodic 2D domain of resolution $\times 64$ 64, with 20 time steps of the vorticity field evolution. We trained models to predict the final 10 time steps from the initial 10, using 1000 training and 200 test samples. Inputs were composed as 3D tensors containing 10 time steps for every spatial point. Padding was introduced only to the time dimension, making the input to be of 64×64×1564$\times$ 64$\times$ 1564 × 64 × 15. We appended spatial coordinates uniformly sampled in $[0,2\pi)$ and a time coordinate with an integer index $t=[0,1,\dots 9]$, which resulted in 4 input channels that were mapped to one channel.
ShapeNet car
This benchmark contains 661 car meshes paired with Reynolds-Averaged Navier–Stokes (RANS) simulations Umetani and Bickel [2018]. We focus on predicting the scalar pressure field at the mesh vertices, which are fixed in size at 3,586 points per mesh. The dataset is split into 561 training and 100 test samples. Target pressures were standardized during training. Following Li et al. [2024], inputs are constructed by sampling a fixed-resolution $64^{3}$ grid over the global bounding box encompassing all meshes. A signed distance function (SDF) is computed per mesh. Grids are padded to $72^{3}$ before being passed to the network.
Ahmed bodies
The Ahmed bodies dataset Ahmed et al. [1984] follows a similar setup to the ShapeNet car but introduces varying inlet velocities as an additional signal. This constant is appended to the SDF input, resulting in two input channels. Meshes vary in size, containing between 90,000 and 200,000 vertices. The grid resolution, padding, and target scaling match those used in the ShapeNet car.
The GINO-decoder architecture Li et al. [2024] used for both ShapeNet and Ahmed datasets requires precomputed fixed-radius neighborhoods to map grid samples to mesh vertices. We use the radii from the original paper: 0.050.050.05 for the ShapeNet car and 0.0350.0350.035 for Ahmed bodies.
For models trained on non-uniform data, we additionally apply Transformer-style positional encoding Vaswani et al. [2017] to the input features, grid locations, and mesh vertices.
Results in the main text are reported on the unscaled targets. Raw dataset specifications are provided in Table 3, and data transformations are detailed in Table 4.
Appendix B Configurations of trained architectures
Table 5: Baseline configurations. GF indicates Gradient Features
Dataset Model $d_{c}$ Modes Rank $r$ GF Extra parameters
Darcy FNO 32 [12, 12] 1 False
flow FNOg 32 [12, 12] 1 True
TFNO 32 [12, 12] $10^{-4}$ False
TFNOg 32 [12, 12] $10^{-4}$ True
[…]
DINOZAUR 32 [12, 12] No True
NS 2D+T FNO 64 [16, 16, 4] 1 False
FNOg 64 [16, 16, 4] 1 True
TFNO 64 [16, 16, 4] $10^{-4}$ False
TFNOg 64 [16, 16, 4] $10^{-4}$ True
[…]
DINOZAUR 64 [16, 16, 4] No True
ShapeNet GINO 64 [24, 24, 24] 0.4 False
car GINOg 64 [24, 24, 24] 0.4 True
TGINO 64 [24, 24, 24] $10^{-4}$ False
TGINOg 64 [24, 24, 24] $10^{-4}$ True
[…]
DINOZAUR 64 [24, 24, 24] No True
GINOD 64 [24, 24, 24] 0.4 False
TGINOD 64 [24, 24, 24] $10^{-4}$ False $\mathrm{D}=0.1$
DINOZAURD 64 [24, 24, 24] No True
GINOL 64 [24, 24, 24] 0.4 False $\sigma_{H}^{2}=e^{-4},$
TGINOL 64 [24, 24, 24] $10^{-4}$ False $C=10^{6},$
DINOZAURL 64 [24, 24, 24] No True $\alpha=500$
DINOZAURB 64 [24, 24, 24] No True $\mathbf{N}_{(\ln\tau)}(\ln 0.01,1),$
$\mathbf{N}_{(\ln\tau|\mathcal{D})}(-5,0.5)$
Ahmed GINO 64 [24, 24, 24] 0.4 False
bodies GINOg 64 [24, 24, 24] 0.4 True
TGINO 64 [24, 24, 24] $10^{-4}$ False
TGINOg 64 [24, 24, 24] $10^{-4}$ True
[…]
DINOZAUR 64 [24, 24, 24] No True
All results reported in our paper are from experiments conducted specifically for this study to ensure a fair and consistent comparison across models. Architectural settings for the baselines are summarized in Table 5. All models featured 4 FNO blocks with GELU activations. The channel width $d_{c}$ and the number of modes were selected to align with the original configurations of FNO Li et al. [2021] and GINO Li et al. [2024], with the exception of Navier-Stokes data. We increased modes and channels to reflect the complexity of this benchmark. For models employing Tucker factorization, the rank $r$ was either adopted directly from the original implementation (GINO) or chosen to match the total number of parameters in DINOZAUR.
For all models incorporating dropout, the probability for a weight element to be zero was set to $\mathrm{D}=0.1$ during training.
For the Laplace approximation, we fit the Hessian of the Negative log-likelihood loss (assuming homoscedastic noise with $\sigma_{H}^{2}=\exp(-4)$) with respect to the parameters $\theta_{\mathcal{Q}}$ of the final (linear) layer $\mathcal{Q}$ of each model. Following Ritter et al. [2018], we scale the Hessian and regularize it by adding a multiple of the identity matrix, yielding
$H_{C,\alpha}=C\mathbb{E}\left[-\frac{\partial^{2}\log p(\mathcal{D}\mid\theta_{\mathcal{Q}})}{\partial\theta^{2}}\right]+\alpha I$
where the expectation is approximated via Monte Carlo samples, and $C$ and $\alpha$ are hyperparameters selected on a validation set. In our experiments, we set $C=10^{6},\,\alpha=500$
DINOZAURB requires us to set the prior distribution parameters for each $\ln\tau_{i}^{c}$ and specify the initial values for the approximate posterior $\ln\tau_{i}^{c}\mid\mathcal{D}$; we initialize this distribution with a full-rank covariance matrix within each NO block. Distribution parameters are informed by the time distributions in the deterministic experiment on ShapeNet car (Figure 5).
Table 6: Baseline optimization parameters
Dataset Model Optimizer N epochs Scheduler LR
Darcy FNO AdamW 150 One-Cycle $10^{-3}$
flow FNOg AdamW 150 One-Cycle $10^{-3}$
TFNO AdamW Schedule-Free 150 None $10^{-3}$
TFNOg AdamW 150 One-Cycle $10^{-3}$
[…]
DINOZAUR AdamW 150 One-Cycle $10^{-3}$
NS 2D+T FNO AdamW 150 One-Cycle $10^{-3}$
FNOg AdamW 150 One-Cycle $10^{-3}$
TFNO AdamW 150 One-Cycle $10^{-3}$
TFNOg AdamW 150 One-Cycle $10^{-3}$
[…]
DINOZAUR AdamW 150 One-Cycle $10^{-2}$
ShapeNet GINO AdamW Schedule-Free 150 None $10^{-3}$
car GINOg AdamW Schedule-Free 150 None $10^{-3}$
TGINO AdamW Schedule-Free 150 None $10^{-3}$
TGINOg AdamW 150 One-Cycle $10^{-3}$
[…]
DINOZAUR AdamW 150 One-Cycle $10^{-3}$
GINOD AdamW Schedule-Free 150 None $10^{-3}$
GINOL AdamW Schedule-Free 150 None $10^{-3}$
TGINOD AdamW Schedule-Free 150 None $10^{-3}$
TGINOL AdamW Schedule-Free 150 None $10^{-3}$
DINOZAURD AdamW 150 One-Cycle $10^{-3}$
DINOZAURL AdamW 150 One-Cycle $10^{-3}$
DINOZAURB AdamW 200 One-Cycle $10^{-3}$
Ahmed GINO AdamW Schedule-Free 150 None $10^{-3}$
bodies GINOg AdamW Schedule-Free 150 None $10^{-3}$
TGINO AdamW 150 One-Cycle $10^{-3}$
TGINOg AdamW 150 One-Cycle $10^{-3}$
[…]
DINOZAUR AdamW 150 One-Cycle $5\times 10^{-3}$
Optimization hyperparameters are detailed in Table 6. All models were trained using either AdamW Schedule-Free Defazio et al. [2024] or AdamW combined with a One-Cycle scheduler Smith and Topin [2019], weight decay set to $10^{-4}$. Both options allow for faster convergence with fewer epochs and reduce sensitivity to hyperparameter choices. The final setup was selected by comparing both options and reporting the results for the configuration that achieved the best performance. A limited learning rate sweep was conducted in the range $[10^{-4},10^{-2}]$, with early termination applied to underperforming runs. For the One-Cycle scheduler, the learning rate in the table refers to the max LR parameter.
Appendix C Distributions of diffusion times
Figure 5: Diffusion times gathered from FNO blocks of deterministic DINOZAUR trained on non-uniform datasets.
Having trained the deterministic DINOZAUR on the non-uniform datasets ShapeNet car and Ahmed bodies, we analyze the behavior of the learned diffusion times across the network. Figure 5 shows the empirical distributions of the 64 diffusion times extracted from each block of the neural operator. We observe two main trends. First, diffusion times generally decrease with network depth. This aligns with the findings of Sharp et al. [2022] and suggests that deeper blocks capture finer-scale features, whereas earlier blocks encode broader, more global structures. Second, diffusion times learned on the Ahmed bodies dataset are consistently larger than those learned on the ShapeNet car. We attribute this to the presence of varying inlet velocities in the Ahmed bodies dataset, which increases the variability in the data and may require coarser diffusion scales to capture the broader range of feature patterns effectively.
Figure 6: Prior distribution set according to $\ln\tau\sim\mathbf{N}(\ln 0.01,1)$ and approximate posteriors gathered from the NO blocks of DINOZAURB trained on ShapeNet car dataset through $\mathcal{L}_{\mathrm{ELBO}}$ maximization.
For the Bayesian version of our model, we initialize all blocks with the same prior distributions for $\ln\tau$ and likewise set identical initial guesses for the corresponding variational posteriors. This initialization allows each block to adapt its uncertainty about diffusion time scales through learning. To visualize this, we draw samples from the approximate posteriors and estimate their empirical location and scale parameters, characterizing the learned diffusion behavior in the probabilistic setting. Figure 6 illustrates these posterior distributions alongside the prior. We find the same consistent pattern: posteriors systematically shift toward shorter diffusion times as depth increases, and their spread becomes tighter. This mirrors the trend observed in the deterministic setting (Figure 5), reinforcing the interpretation that deeper blocks specialize in capturing high-frequency, localized features. The prior has a strong influence on the resulting posteriors, forcing diffusion times to be lower (closer to the prior mean, which was chosen to promote shorter ranges for message passing).
Figure 7: Lower diagonal full-rank learned covariance matrices of the posterior distributions. Each matrix represents a $d_{c}\times d_{c}$ ($\times 64$ 64) covariance matrix of the multi-dimensional distribution of diffusion times within each block.
Finally, in Figure 7, we provide the full covariance matrices for each of the posteriors that were learned by the network. Analyzing covariances, we note that they are dominated by the diagonals with few spots of strong correlation off-diagonal. This indicates a good separation of information within features.
Appendix D Metrics definitions
Given a test set element index $n=1,\dots,N_{\mathrm{test}}$ and a point evaluation index $j=1,\dots,J_{n}$, let $u_{nj}$ denote the ground truth observation of the target function $u$ for element $n$ at point $j$, and let $\widehat{u}_{nj}$ denote the corresponding neural operator prediction. For baseline models that support uncertainty quantification via sampling, $\widehat{u}_{nj}$ represents the empirical mean over samples, and $\widehat{\sigma}_{nj}$ denotes the empirical standard deviation. In all probabilistic experiments, we generated 100 samples per prediction.
Relative L2
The RL2 loss is a normalized metric that quantifies the discrepancy between predicted outputs and ground truth values, scaled by the squared norm of the ground truth. This normalization allows for fair comparisons across datasets or tasks with varying magnitudes and units.
$\mathrm{RL}_{2}=\frac{1}{N_{\mathrm{test}}}\sum_{n=1}^{N_{\mathrm{test}}}\frac{\sqrt{\sum_{j=1}^{J_{n}}(u_{nj}-\widehat{u}_{nj})^{2}}}{\sqrt{\sum_{j=1}^{J_{n}}u_{nj}^{2}}}.$
Negative log-likelihood
NLL reflects the mean log-likelihood of the true values under the predicted Gaussian distribution. Lower NLL indicates better uncertainty calibration and fit to the data.
$\mathrm{NLL}=\frac{1}{N_{\mathrm{test}}}\sum_{n=1}^{N_{\mathrm{test}}}\frac{1}{J_{n}}\sum_{j=1}^{J_{n}}\left[\ln(\widehat{\sigma}_{nj}\sqrt{2\pi})+\frac{(u_{nj}-\widehat{u}_{nj})^{2}}{2\widehat{\sigma}_{nj}^{2}}\right].$
This averaged formula was used only for reporting metrics for probabilistic models.
Miscalibration area
The MA is a scalar metric that quantifies the discrepancy between predicted confidence levels and the actual observed frequencies of events, providing a measure of uncertainty calibration. It captures the area between the model’s empirical calibration curve and the ideal diagonal (perfect calibration). Lower values indicate better uncertainty calibration. Let $\pi^{(k)}=\frac{k}{K}$ for $k=0,\dots,K$, denote the expected coverage levels. For each test sample $n$, the observed coverage at level $\pi^{(k)}$ is given by:
$\widehat{\pi}_{n}^{(k)}=\frac{1}{J_{n}}\sum_{j=1}^{J_{n}}\mathbbm{1}\left(\frac{|u_{nj}-\widehat{u}_{nj}|}{\widehat{\sigma}_{nj}}\leq\Phi^{-1}\left(\frac{1+\pi^{(k)}}{2}\right)\right),$
where $\Phi^{-1}$ is the inverse of the standard normal cumulative distribution function. Then, the miscalibration area across all elements of the test set is given by averaging the area estimates for each of the elements:
$\mathrm{MA}=\frac{1}{N_{\mathrm{test}}}\sum_{n=1}^{N_{\mathrm{test}}}\sum_{k=0}^{K-1}\left|\frac{(\pi^{(k+1)}-\pi^{(k)})}{2}\left[\widehat{\pi}_{n}^{(k)}-\pi^{(k)}+\widehat{\pi}_{n}^{(k+1)}-\pi^{(k+1)}\right]\right|.$
Interval score
IS is another metric for evaluating predictive uncertainty. It balances the width of the prediction interval with a penalty for ground truth values falling outside the interval. For a chosen coverage level $p\in(0,1)$, define the lower and upper predictive bounds:
$\underline{b_{nj}^{(p)}}=\widehat{u}_{nj}+\widehat{\sigma}_{nj}\cdot\Phi^{-1}\left(\frac{1-p}{2}\right)\quad\text{and}\quad\overline{b_{nj}^{(p)}}=\widehat{u}_{nj}+\widehat{\sigma}_{nj}\cdot\Phi^{-1}\left(\frac{1+p}{2}\right)$
where $\Phi^{-1}$ is the inverse of the standard normal cumulative distribution function. Then the interval score at level $p$ for prediction at point $j$ in sample $n$ is:
$\mathrm{IS}_{nj}^{(p)}=\left(\overline{b_{nj}^{(p)}}-\underline{b_{nj}^{(p)}}\right)+\frac{2}{1-p}\cdot\left(\underline{b_{i}^{(p)}}-u_{i}\right)\cdot\mathbbm{1}\left(u_{i}<\underline{b_{i}^{(p)}}\right)+\frac{2}{1-p}\cdot\left(u_{i}-\overline{b_{i}^{(p)}}\right)\cdot\mathbbm{1}\left(u_{i}>\overline{b_{i}^{(p)}}\right).$
To compute the overall interval score, we average over a predefined set of $P$ equally spaced coverage levels ($p=[0.01,0.02,\dots 0.99]$):
$\mathrm{IS}=\frac{1}{N_{\mathrm{test}}}\sum_{n=1}^{N_{\mathrm{test}}}\frac{1}{P}\sum_{p}\frac{1}{J_{n}}\sum_{j=1}^{J_{n}}\mathrm{IS}_{nj}^{(p)}.$
Appendix E Elaboration on the universal approximation
Here, we formulate the statement of Proposition 1, reported in the main text, more strictly.
Following the exposition of Lanthaler et al. [2023], let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary and let $\mathcal{A}(\Omega,\mathbb{R}^{d_{a}})$, $\mathcal{U}(\Omega,\mathbb{R}^{d_{u}})$, $\mathcal{V}(\Omega,\mathbb{R}^{d_{c}})$ denote Banach space of functions on $\Omega$. DINOZAUR architecture defines a mapping
$\mathcal{N}_{\theta,\tau}:\mathcal{A}(\Omega,\mathbb{R}^{d_{a}})\rightarrow\mathcal{U}(\Omega,\mathbb{R}^{d_{u}}),$
which can be written as a composition of a form:
$\mathcal{N}_{\theta,\tau}=\mathcal{Q}\circ\mathcal{N}_{\theta_{M},\tau_{M}}\circ\dots\circ\mathcal{N}_{\theta_{1},\tau_{1}}\circ\mathcal{P},$
consisting of lifting layer $\mathcal{P}$, hidden blocks $\mathcal{N}_{\theta_{i},\tau_{i}},\,i=1,\dots,M$, and a projection layer $\mathcal{Q}$. Given channel dimension $d_{c}$, the lifting layer $\mathcal{P}$ is given by mapping:
$\mathcal{P}:\mathcal{A}(\Omega,\mathbb{R}^{d_{a}})\rightarrow\mathcal{V}(\Omega,\mathbb{R}^{d_{c}}),\quad a(x)\mapsto P(a(x),x),$
where $P:\mathbb{R}^{d_{a}}\times\Omega\rightarrow\mathbb{R}^{d_{c}}$ is a learnable neural network acting between finite-dimensional Euclidean spaces. For $i=1,\dots,M$, each block $\mathcal{N}_{\theta_{i},\tau_{i}}$ is of the form
$\mathcal{N}_{\theta_{i},\tau_{i}}[v](x):=\sigma\left(W_{i}^{\mathrm{skip}}v(x)+b_{i}+W_{i}^{\mathrm{mix}}\begin{bmatrix}\mathcal{F}^{-1}\bigl{[}\exp(-4\pi^{2}\|k\|^{2}\tau_{i})\odot\mathcal{F}[v](k)\bigr{]}(x)\\ \mathcal{G}_{W_{i}^{\mathrm{grad}}}\Bigl{[}\mathcal{F}^{-1}\bigl{[}\exp(-4\pi^{2}\|k\|^{2}\tau_{i})\odot\mathcal{F}[v](k)\bigr{]}\Bigr{]}(x)\end{bmatrix}\right).$
Each hidden block defines a mapping $\mathcal{N}_{\theta_{i},\tau_{i}}:\mathcal{V}(\Omega,\mathbb{R}^{d_{c}})\rightarrow\mathcal{V}(\Omega,\mathbb{R}^{d_{c}})$. For $i=1,\dots,M$, the matrices $W_{i}^{\mathrm{skip}},W_{i}^{\mathrm{grad}}\in\mathbb{R}^{d_{c}\times d_{c}}$, $W_{i}^{\mathrm{mix}}\in\mathbb{R}^{d_{c}\times 2d_{c}}$ diffusion times $\tau_{i}\in\mathbb{R}^{d_{c}}$ and bias $b_{i}\in\mathbb{R}^{d_{c}}$ are learnable parameters. The non-linearity $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ acts element-wise and is assumed to be smooth, non-polynomial, and Lipschitz-continuous. Finally, the projection layer $\mathcal{Q}$ is given by a mapping:
$\mathcal{Q}:\mathcal{V}(\Omega,\mathbb{R}^{d_{c}})\rightarrow\mathcal{U}(\Omega,\mathbb{R}^{d_{u}}),\quad v(x)\mapsto Q(v(x),x),$
where $Q:\mathbb{R}^{d_{c}}\times\Omega\rightarrow\mathbb{R}^{d_{u}}$ is also a learnable neural network acting between finite-dimensional Euclidean spaces.
From now on, we will focus on the case of a single hidden block:
$\mathcal{N}_{\theta,\tau}[a](x)=(\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\tau}\circ\mathcal{P})[a](x).$
Now, we re-state the Proposition 1:
Proposition 1.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary and such that the closure $\overline{\Omega}\subset(0,2\pi)^{d}$. For given integers $s,s^{\prime}$ > 0, let $\mathcal{N}:C^{s}(\overline{\Omega};\mathbb{R}^{d_{a}})\rightarrow C^{s^{\prime}}(\overline{\Omega};\mathbb{R}^{d_{u}})$ be a continuous operator, and fix a compact set $\mathbf{A}\subset C^{s}(\overline{\Omega};\mathbb{R}^{d_{a}})$. Then there exists a continuous, linear operator $\mathcal{E}:C^{s}(\Omega;\mathbb{R}^{d_{a}})\rightarrow C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})$ such that $\mathcal{E}[a]|_{\Omega}=a$ for all $a\in C^{s}(\Omega;\mathbb{R}^{d_{a}})$. Furthermore, for any $\varepsilon>0$, there exists a DINOZAUR such that
$\sup_{a\in\mathbf{A}}\|\mathcal{N}[a]-(\mathcal{N}_{\theta,\tau}\circ\mathcal{E}[a])|_{\Omega}\|_{C^{s^{\prime}}(\overline{\Omega};\mathbb{R}^{d_{u}})}\leq\varepsilon,$
where $\mathcal{N}_{\theta,\tau}=\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\tau}\circ\mathcal{P}:C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})\rightarrow C^{s^{\prime}}(\mathbb{T}^{d},\mathbb{R}^{d_{u}})$.
Proof.
First, we consider a periodic torus $\mathbb{T}^{d}$. Without loss of generality, since the weights of the network are free to be updated, we let the matrices $W^{\mathrm{mix}}_{i}$ be of the block structure
$W^{\mathrm{mix}}_{i}=\begin{bmatrix}\frac{1}{|\mathbb{T}^{d}|}I\\ 0\end{bmatrix},$
with $I\in\mathbb{R}^{d_{c}\times d_{c}}$ being the identity matrix and $0\in\mathbb{R}^{d_{c}\times d_{c}}$ a matrix of all zeros, which renders a hidden block to:
$\mathcal{N}_{\theta_{1},\tau}[v](x)=\sigma\left(W_{1}^{\mathrm{skip}}v(x)+b_{1}+\frac{1}{|\mathbb{T}^{d}|}\mathcal{F}^{-1}\bigl{[}\exp(-4\pi^{2}\|k\|^{2}\tau)\odot\mathcal{F}[v](k)\bigr{]}(x)\right).$
By letting diffusion times $\tau\rightarrow\infty$, we get
$\exp(-4\pi^{2}\|k\|^{2}\tau)$ $=1,\,\|k\|=0,$
$\exp(-4\pi^{2}\|k\|^{2}\tau)$ $\rightarrow 0,\,\|k\|>0,$
which, in the limit, results in the block:
$\mathcal{N}_{\theta_{1},\infty}[v](x)$ $=\sigma\left(W_{1}^{\mathrm{skip}}v(x)+b_{1}+\frac{1}{|\mathbb{T}^{d}|}\mathcal{F}^{-1}\bigl{[}\mathcal{F}[v](0)\bigr{]}(x)\right)$
$=\sigma\left(W_{1}^{\mathrm{skip}}v(x)+b_{1}+\fint_{\mathbb{T}^{d}}v(y)\mathrm{d}y\right),$
making a full network $\mathcal{N}_{\theta,\infty}=\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\infty}\circ\mathcal{P}$ an averaging neural operator in the sense of Equation 2.3 in Lanthaler et al. [2023] on the torus $\mathbb{T}^{d}$. By Corollary 3.3 Lanthaler et al. [2023], $\mathcal{N}_{\theta,\infty}$ is a universal approximator on the torus. Here, we note that the lifting and projection operators defined in Lanthaler et al. [2023] have an explicit positional dependency, and thus, one should understand them as periodized MLPs. 111 Note that Theorem 2.1 in Lanthaler et al. [2023] as it is formulated, applies to bounded Lipschitz domains. Thus, the proof of Corollary 3.3 Lanthaler et al. [2023] is not immediate. One must either follow the proof of Theorem 2.1 to see that it applies to the torus, or first formulate it on the domain $(0,2\pi)^{d}$, invoke the Theorem 2.1, and then argue that it holds when one restricts to periodic functions on $(0,2\pi)^{d}$.
Now consider a general domain $\Omega\subset(0,2\pi)^{d}$. By Lemma 41 by Kovachki et al. [2021] and Lemma A1 in Lanthaler et al. [2023], there exists a continuous, linear extension operator $\mathcal{E}:C^{s}(\Omega;\mathbb{R}^{d_{a}})\rightarrow C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})$ such that $\mathcal{E}[a]|_{\Omega}=a$ and $\mathcal{E}[a]$ is periodic on $[0,2\pi]^{d}$ for all $a\in C^{s}(\Omega;\mathbb{R}^{d_{a}})$. We note that the conditions of the Theorem 9 in Kovachki et al. [2021] are satisfied. By following the proof of this theorem, we apply the extension operator $\mathcal{E}$ and use the universal approximation on the torus, establishing universal approximation on $\Omega$. ∎

### v2

@@
 Our method achieves competitive or superior performance across several PDE benchmarks while providing efficient uncertainty quantification.
+The code is available at https://github.com/PhysicsXLtd/DINOZAUR.
 1 Introduction
@@
 Numerous problems in science and engineering require formulating and solving PDEs, often presenting significant computational and methodological challenges to researchers and practitioners.
-A common strategy for solving PDEs is via conventional numerical methods, such as finite difference, finite element, or finite volume method Larsson and Thomée [2003].
-While accurate, they become prohibitively expensive in complex engineering tasks, especially where rapid iteration or real-time response is critical, such as design optimization or control Martins and Ning [2021].
+A common strategy for solving PDEs is via conventional numerical methods, such as finite difference, finite element, or finite volume method [1].
+While accurate, they become prohibitively expensive in complex engineering tasks, especially where rapid iteration or real-time response is critical, such as design optimization or control [2].
 In recent years, there has been a growing trend toward replacing computationally intensive numerical solvers with neural network-based surrogate models trained to approximate PDE solutions.
-Since discretized PDE domains are commonly represented using point clouds, meshes, or graphs, geometric deep learning Bronstein et al.
-[2021] has emerged as a promising paradigm for this class of problems.
+Since discretized PDE domains are commonly represented using point clouds, meshes, or graphs, geometric deep learning [3] has emerged as a promising paradigm for this class of problems.
 In particular, message passing layers, initially developed for geometric and relational data, have proven effective due to their structural resemblance to physical convolution operators.
-These layers are especially suitable for capturing the spatial dependencies inherent in PDE systems, making them well-suited for learning dynamics over irregular or non-Euclidean domains Pfaff et al.
-[2020], Brandstetter et al.
-[2022].
+These layers are especially suitable for capturing the spatial dependencies inherent in PDE systems, making them well-suited for learning dynamics over irregular or non-Euclidean domains [4, 5].
 However, geometric deep learning methods often lack key properties required for solving PDEs, motivating alternative approaches.
 A PDE – its parameters, equation structure, and boundary conditions – defines an operator that maps input functions to solution functions.
-Neural operators (NOs) Lu et al.
-[2019], Kovachki et al.
-[2023, 2024] have recently emerged as a leading framework for approximating such solution operators and, more generally, for tackling the operator learning problem implicitly present in PDEs.
-These models incorporate inductive biases that improve performance and offer discretization invariance and universal approximation Kovachki et al.
-[2023], making them a strong alternative to geometric deep learning.
-A prominent example is the Fourier Neural Operator (FNO) Li et al.
-[2021], which learns transformations in the spectral domain via the Fast Fourier Transform.
-While FNOs are restrictive in requiring input data on a regular grid, they remain a widely used backbone in neural operator architectures Li et al.
-[2024].
-Despite their success on problems such as Darcy flow and Burgers’ equation, FNOs are known to be overparameterized Tran et al.
-[2023], Kossaifi et al.
-[2024].
+Neural operators (NOs) [6, 7, 8] have recently emerged as a leading framework for approximating such solution operators and, more generally, for tackling the operator learning problem implicitly present in PDEs.
+These models incorporate inductive biases that improve performance and offer discretization invariance and universal approximation [7], making them a strong alternative to geometric deep learning.
+A prominent example is the Fourier Neural Operator (FNO) [9], which learns transformations in the spectral domain via the Fast Fourier Transform.
+While FNOs are restrictive in requiring input data on a regular grid, they remain a widely used backbone in neural operator architectures [10].
+Despite their success on problems such as Darcy flow and Burgers’ equation, FNOs are known to be overparameterized [11, 12].
 Their design leads to exponential growth with domain dimensionality, causing overfitting and significant memory overhead.
@@
 It enhances model outputs by accounting for different sources of error, helping guide more informed and reliable downstream use.
-Uncertainty modeling has long relied on Bayesian inference Mackay [1992], ensembles Dong et al.
-[2020], and conformal prediction Shafer and Vovk [2008] to produce confidence estimates.
-While some of these methods have been adapted to deep learning with varying success Gal and Ghahramani [2016], Zhang et al.
-[2021], Jospin et al.
-[2022], UQ for GNNs and neural operators Psaros et al.
-[2023] remains underdeveloped, often relying on post hoc methods ill-suited to their structure.
+Uncertainty modeling has long relied on Bayesian inference [13], ensembles [14], and conformal prediction [15] to produce confidence estimates.
+While some of these methods have been adapted to deep learning with varying success [16, 17, 18], UQ for GNNs and neural operators [19] remains underdeveloped, often relying on post hoc methods ill-suited to their structure.
 To address the scalability issues and limited UQ methods in FNOs, we present DINOZAUR – a diffusion-based neural operator parametrization with uncertainty quantification.
 Our diffusion multiplier is inspired by the heat kernel in the Fourier basis and requires only one learnable parameter per channel, resulting in superior scalability without sacrificing predictive performance.
-This physically motivated choice of multiplier allows us to place meaningful probabilistic priors on the learnable time parameters, recover the posterior via inexpensive variational inference (VI) Blei et al.
-[2017] of these parameters, and efficiently sample spatially correlated functions.
+This physically motivated choice of multiplier allows us to place meaningful probabilistic priors on the learnable time parameters, recover the posterior via inexpensive variational inference (VI) [20] of these parameters, and efficiently sample spatially correlated functions.
 We summarize our contributions as follows:
@@
 2 Related work
-The area of designing learnable approximation for PDE solutions is vast and rapidly growing Wang et al.
-[2024], Huang et al.
-[2025].
+The area of designing learnable approximation for PDE solutions is vast and rapidly growing [21, 22].
 Given the increasing availability of data, the need to incorporate inductive biases, and the inherently geometric structure of many physical problems, two classes of deep learning models have come to dominate the field: geometric deep learning models and neural operators.
 Geometric deep learning
-Graph neural networks (GNNs) Scarselli et al.
-[2008] offer a flexible framework for learning on non-Euclidean domains.
-Many GNNs implement spatial convolution by performing learnable aggregation of local neighborhoods Niepert et al.
-[2016], Chamberlain et al.
-[2021].
-Neural PDE solvers increasingly rely on spatial GNN layers to model and predict physical phenomena across various domains, including control Li et al.
-[2019], deformable material simulation Sanchez-Gonzalez et al.
-[2020], and spatiotemporal mesh evolution Pfaff et al.
-[2020].
-Specific architectures target building robust autoregressive solvers Brandstetter et al.
-[2022] and encoding the geometric boundary conditions Mayr et al.
-[2023].
+Graph neural networks (GNNs) [23] offer a flexible framework for learning on non-Euclidean domains.
+Many GNNs implement spatial convolution by performing learnable aggregation of local neighborhoods [24, 25].
+Neural PDE solvers increasingly rely on spatial GNN layers to model and predict physical phenomena across various domains, including control [26], deformable material simulation [27], and spatiotemporal mesh evolution [4].
+Specific architectures target building robust autoregressive solvers [5] and encoding the geometric boundary conditions [28].
 However, their performance often suffers from sensitivity to discretization, limiting generalization across mesh resolutions.
-Spectral GNNs define convolutions using the graph Laplacian’s eigen-decomposition Bruna et al.
-[2014], Kipf and Welling [2017], often with polynomial approximations of spectral filters Defferrard et al.
-[2016], Levie et al.
-[2018].
-These have seen success in 3D shape analysis, particularly with functional maps Ovsjanikov et al.
-[2012].
-DiffusionNet Sharp et al.
-[2022] exemplifies this approach by using the spectral form of the heat kernel for a simple and interpretable kernel parametrization.
+Spectral GNNs define convolutions using the graph Laplacian’s eigen-decomposition [29, 30], often with polynomial approximations of spectral filters [31, 32].
+These have seen success in 3D shape analysis, particularly with functional maps [33].
+DiffusionNet [34] exemplifies this approach by using the spectral form of the heat kernel for a simple and interpretable kernel parametrization.
 It exhibits strong performance and generalization; we note that such a kernel is applicable in arbitrary domains.
 Neural operators
-Neural operators are a powerful class of models for learning mappings between infinite-dimensional spaces in a resolution-invariant manner Kovachki et al.
-[2023], offering an alternative to GNNs.
+Neural operators are a powerful class of models for learning mappings between infinite-dimensional spaces in a resolution-invariant manner [7], offering an alternative to GNNs.
 Lu et al.
 [6] introduced DeepONet, inspired by operator learning and universal approximation theory, using a branch-trunk architecture to encode inputs and evaluate the solution at arbitrary locations.
-A parallel line of work introduced the Graph Neural Operator (GNO) architecture Anandkumar et al.
-[2020], Li et al.
-[2020], transferring the message passing intuition from GNNs to NOs.
-A family of neural operators decomposes the input function into a basis defined by one of the popular transforms: Fourier Li et al.
-[2021], Laplace Cao et al.
-[2024], Wavelet Tripura and Chakraborty [2023].
-In particular, FNOs are popular for their performance and interpretability despite needing a regular grid sampling, and are commonly used in other networks, such as Geo-FNO Li et al.
-[2023], U-FNO Wen et al.
-[2022], and Geometry-Informed Neural Operator (GINO) Li et al.
-[2024], to name a few.
-However, FNOs face scalability and generalization issues due to exponential parameter growth with domain dimensionality Tran et al.
-[2023].
-Efforts to improve this, such as FFT factorization Tran et al.
-[2023], low-rank tensor decomposition Kossaifi et al.
-[2024], and MLPs in the spectral domain Xiao et al.
-[2024], have reduced growth but still rely on spatial dimensions.
+A parallel line of work introduced the Graph Neural Operator (GNO) architecture [35, 36], transferring the message passing intuition from GNNs to NOs.
+A family of neural operators decomposes the input function into a basis defined by one of the popular transforms: Fourier [9], Laplace [37], Wavelet [38].
+In particular, FNOs are popular for their performance and interpretability despite needing a regular grid sampling, and are commonly used in other networks, such as Geo-FNO [39], U-FNO [40], and Geometry-Informed Neural Operator (GINO) [10], to name a few.
+However, FNOs face scalability and generalization issues due to exponential parameter growth with domain dimensionality [11].
+Efforts to improve this, such as FFT factorization [11], low-rank tensor decomposition [12], and MLPs in the spectral domain [41], have reduced growth but still rely on spatial dimensions.
 Functional uncertainty quantification
-UQ in deep learning has been widely studied, with popular methods including MC Dropout Gal and Ghahramani [2016], Laplace approximation MacKay [1992], ensembles Rahaman et al.
-[2021], and non-parametric conformal prediction Zhang et al.
-[2021], which typically require extra computations or careful calibration.
-In the context of functional UQ – crucial for neural operators – the landscape is similarly limited.
-Recent approaches include applying a post hoc Laplace approximation to the final layer of the NO Magnani et al.
-[2022], using conformal methods after training Ma et al.
-[2024], or learning an uncertainty measure as an auxiliary output Bülte et al.
-[2025].
-However, none of these strategies model uncertainty within the NO; instead, they adopt generic methods for neural networks and neglect the inductive biases of neural operators.
+UQ in deep learning has been widely studied, with popular methods including MC Dropout [16], Laplace approximation [42], ensembles [43], and non-parametric conformal prediction [17], which typically require extra computations or careful calibration.
+In the context of functional UQ – crucial for neural operators – the landscape is further limited.
+Recent approaches include applying a post hoc Laplace approximation to the final layer of the NO [44], using conformal methods after training [45], or learning an uncertainty measure as an auxiliary output [46].
+Another family of approaches leverages diffusion-based generative models for uncertainty quantification in physics surrogates [47, 48, 49, 50].
+While these methods use diffusion models to capture statistics of high-dimensional dynamical systems and generate realistic samples, they require solving a differential equation for reverse diffusion numerically through multiple denoising steps, which is computationally intensive compared to deterministic forward passes in neural operators.
+Despite these various approaches, none of them model uncertainty within the NO itself; instead, they adopt generic methods for neural networks and neglect the inductive biases of neural operators.
 3 Diffusion multiplier parametrization for neural operators
@@
 where $f(x)$ is a fixed forcing function and $g(x)$ is a boundary condition.
-This formulation is general enough to include varying geometric boundaries by including the signed distance function of a geometry in functions from $\mathcal{A}$ and defining a solution as an extension to the whole domain Stein [1970], Li et al.
-[2024].
+This formulation is general enough to include varying geometric boundaries by including the signed distance function of a geometry in functions from $\mathcal{A}$ and defining a solution as an extension to the whole domain [51, 10].
 We assume the PDE induces a continuous solution map $\mathcal{N}:\mathcal{A}\rightarrow\mathcal{U}$ that maps inputs $a$ to solutions $u$, and we are seeking to find its parametrized approximation $\mathcal{N}_{\theta}$ with $p$ parameters.
@@
 $v_{i+1}(x)=\mathcal{N}_{\theta_{i}}[v_{i}](x)=\sigma\left(W^{\mathrm{skip}}_{i}v_{i}(x)+b_{i}+\int_{\Omega}\kappa_{\gamma_{i}}(x,y)v_{i}(y)\mathrm{d}y\right),\,i=1,\dots,M,$ (3)
-where $\kappa_{\gamma_{i}}$ is a kernel, matrix $W^{\mathrm{skip}}\in\mathbb{R}^{d_{c}\times d_{c}}$, $b_{i}\in\mathbf{R}^{d_{c}}$ is a bias and $d_{c}$ is the number of channels.
+where $\kappa_{\gamma_{i}}$ is a kernel, matrix $W^{\mathrm{skip}}\in\mathbb{R}^{d_{c}\times d_{c}}$, $b_{i}\in\mathbb{R}^{d_{c}}$ is a bias and $d_{c}$ is the number of channels.
 FNO restricts domain to be periodic torus $\Omega=\mathbb{T}^{d}$, kernels to be stationary and represents the integral transform $\mathcal{I}$ in the following way:
@@
 Diffusion multiplier
-Motivated by the success of DiffusionNet Sharp et al.
-[2022] and its resemblance to FNO, we choose to restrict the form of $\kappa_{\gamma}$ to that of the heat kernel for our architecture.
+Motivated by the success of DiffusionNet [34] and its resemblance to FNO, we choose to restrict the form of $\kappa_{\gamma}$ to that of the heat kernel for our architecture.
 Consider the heat equation on the torus $\Omega=\mathbb{T}^{d}$:
@@
 $v(0,x)$ $=v_{0}(x),$
-with a solution Strauss [2007] given by:
+with a solution [52] given by:
 $v(\tau,x)=\mathcal{I}_{\tau}[v_{0}](x)=\mathcal{F}^{-1}\bigl[R_{\tau}(k)\mathcal{F}[v_{0}](k)\bigr](x)=\mathcal{F}^{-1}\bigl[\exp(-\lambda_{k}\tau)\odot\mathcal{F}[v_{0}](k)\bigr](x),$ (6)
@@
 The final block architecture is:
-$v_{i+1}(x)=\sigma\left(W^{\mathrm{skip}}_{i}v_{i}(x)+b_{i}+W_{i}^{\mathrm{mix}}\begin{bmatrix}\mathcal{I}_{\tau_{i}}[v_{i}](x)\\ \big{(}\mathcal{G}_{\theta_{i}}\circ\mathcal{I}_{\tau_{i}}\big{)}[v_{i}](x)\end{bmatrix}\right).$ (8)
+[…]
 Substituting the original blocks in (3) in the neural operator $\mathcal{N}_{\theta}$ with modified blocks from (8) results in DINOZAUR architecture that we denote as $\mathcal{N}_{\theta,\tau}$, with all non-diffusion parameters lumped together in $\theta$ and separated from $\tau$.
@@
 3.3 Bayesian inference for diffusion multiplier
-$a$$u$$\tau$$\sigma$$\theta$ $N$
+[…]
 Figure 2: (a) Probabilistic graph for DINOZAUR, showing how the latent diffusion-times, $\tau$, relate to the observations and non-mechanistic model parameters, $\theta,\sigma$.
@@
 See Figure 2(a) for the corresponding graphical model.
-We also collect training data in a set of discrete observations, $\mathcal{D}=\left\{\{a_{nk}\}_{k=1}^{K_{n}},\{u_{nj}\}_{j=1}^{J_{n}}\right\}_{n=1}^{N}$, with possibly different sets of evaluation points $\{x_{k}\}$, $\{y_{j}\}$.
+We also collect training data in a set of discrete observations, $\mathcal{D}=\left\{\{a_{nk}\}_{k=1}^{K_{n}},\{u_{nj}\}_{j=1}^{J_{n}}\right\}_{n=1}^{N}$, with possibly different sets of function evaluation points at the input and output.
 We seek to recover the approximate posterior for the diffusion time, $p_{\theta,\sigma}(\tau|\mathcal{D})$, according to (9), which yields a posterior for the diffusion multipliers, $R_{\tau}$, and the entire neural operator $\mathcal{N}_{\theta,\tau}$.
@@
 Due to their parsimonious and mechanistic role in controlling message propagation, diffusion time parameters avoid the redundancy and symmetries that challenge VI for typical neural network weights, allowing for better identifiability and justifying a full-rank approximation within each block.
-We then seek to find $\phi$ that minimizes the Kullback-Leibler divergence (KL) from the $q(\tau;\phi)$ to the true posterior.
-As this objective is still intractable Murphy [2022], we instead maximize a lower bound (ELBO):
+We then seek to find $\phi$ that minimizes the Kullback-Leibler divergence (KL) from $q(\tau;\phi)$ to the true posterior.
+As this objective is still intractable [53], we instead maximize a lower bound (ELBO):
 $\mathcal{L}_{\rm ELBO}(\theta,\sigma;\phi)=$ $\mathbb{E}_{q(\tau;\phi)}[\ln p_{\theta,\sigma}(\mathcal{D}|\tau)]-D_{\rm KL}[q(\tau;\phi)\|p(\tau)],$ (12)
-which allows to learn jointly the network parameters $(\theta,\sigma)$ and variational parameters $\phi$.
-During inference, we replace the true posterior $p_{\theta,\sigma}$ with its approximation $q(\tau;\phi)$ in (10).
+which allows to jointly learn the network parameters $(\theta,\sigma)$ and variational parameters $\phi$.
+To predict for a new instance, we replace the true posterior $p_{\theta,\sigma}$ with its approximation $q(\tau;\phi)$ in (10).
 4 Experiments
 Table 1: Comparison of DINOZAUR against the baselines in deterministic setting, reporting the $\mathrm{RL}_{2}$ values on test split.
+Mean and standard deviation values are computed from three trained models.
 The best result is bold, second best is $\underline{\text{underlined}}$
-Model # params $\mathrm{RL}_{2}\downarrow$ params $\mathrm{RL}_{2}\downarrow$
-(×$(\times 10^{3})$ (×$(\times 10^{-2})$ (×$(\times 10^{3})$ (×$(\times 10^{-2})$
+Model # params $\mathrm{RL}_{2}\downarrow$ # params $\mathrm{RL}_{2}\downarrow$
+$(\times 10^{3})$ $(\times 10^{-2})$ $(\times 10^{3})$ $(\times 10^{-2})$
 Darcy flow Navier-Stokes
-FNO 361.4 $\underline{0.64}$ 12,650.8 $\underline{15.96}$
-FNOg 371.9 0.86 12,692.2 17.24
-TFNO 17.7 4.44 69.1 21.39
-TFNOg 28.2 2.28 110.6 19.40
+FNO 361.4 $\underline{0.90\pm 0.03}$ 12,650.8 $\underline{15.90\pm 0.20}$
+FNOg 371.9 $0.91\pm 0.07$ 12,692.2 $17.08\pm 0.12$
+TFNO 17.7 $8.04\pm 0.86$ 69.1 $21.48\pm 0.27$
+TFNOg 28.2 $4.16\pm 0.61$ 110.6 $19.47\pm 0.58$
 […]
-DINOZAUR (Ours) 28.0 0.39 109.7 15.86
+DINOZAUR (Ours) 28.0 $\mathbf{0.75\pm 0.10}$ 109.7 $\mathbf{15.68\pm 0.21}$
 ShapeNet car Ahmed bodies
-GINO 49,608.6 8.30 49,612.9 8.51
-GINOg 49,650.1 7.03 49,645.3 $\underline{8.07}$
-TGINO 149.8 7.42 154.0 8.13
-TGINOg 191.3 8.09 195.5 8.86
+GINO 49,608.6 $8.03\pm 0.39$ 49,612.9 $8.57\pm 0.13$
+GINOg 49,650.1 $7.76\pm 0.64$ 49,645.3 $8.25\pm 0.16$
+TGINO 149.8 $\underline{7.27\pm 0.16}$ 154.0 $\underline{8.19\pm 0.49}$
+TGINOg 191.3 $8.30\pm 0.38$ 195.5 $8.23\pm 0.64$
 […]
-DINOZAUR (Ours) 180.4 $\underline{7.08}$ 184.6 7.96
+DINOZAUR (Ours) 180.4 $\mathbf{7.04\pm 0.06}$ 184.6 $\mathbf{7.99\pm 0.07}$
 We evaluate the performance of the DINOZAUR architecture on several standard benchmarks.
@@
 Ablation studies and computational performance comparisons are detailed in Section 4.3.
-All models are implemented in PyTorch Paszke et al.
-[2017] and trained on a single NVIDIA A100 GPU for 150 epochs with AdamW Loshchilov and Hutter [2019] and One-Cycle scheduler Smith and Topin [2019] or AdamW Schedule-Free Defazio et al.
-[2024] with learning rate (in most cases) set to $10^{-3}$.
+All models are implemented in PyTorch [54] and trained on a single NVIDIA A100 GPU for 150 epochs with AdamW [55] and One-Cycle scheduler [56] or AdamW Schedule-Free [57] with learning rate (in most cases) set to $10^{-3}$.
 See the Appendix B for details.
+All experiments were repeated three times with random initialization of trainable parameters.
+Reported results include the mean and standard deviation over all runs.
 Benchmarks
-We evaluate our method on two regularly sampled datasets from Li et al.
-[2021] and two datasets with irregular domains used in Li et al.
-[2024].
+We evaluate our method on two regularly sampled datasets from [9] and two datasets with irregular domains used in [10].
 For the 2D Darcy flow, we learn a mapping from a scalar permeability field of resolution $241^{2}$ to a scalar pressure field, applying standard scaling to both.
@@
 We learn to map the first 10 time steps of vorticity to the last 10, without applying data scaling.
-For further dataset and PDE details, we refer to Takamoto et al.
-[2022].
-The irregular-domain benchmarks include two datasets for predicting surface pressure on 3D meshes: ShapeNet car Umetani and Bickel [2018] and Ahmed bodies Ahmed et al.
-[1984].
+For further dataset and PDE details, we refer to [58].
+The irregular-domain benchmarks include two datasets for predicting surface pressure on 3D meshes: ShapeNet car [59] and Ahmed bodies [60].
 Following the protocol of Li et al.
@@
 In contrast, DINOZAUR maintains strong results across domains, demonstrating the effectiveness of diffusion-based message passing as a parameter-efficient alternative.
-Table 2: Evaluation of the uncertainty output quality of DINOZAUR and baseline models. The best result is bold, second best is $\underline{\text{underlined}}$
+Table 2: Evaluation of the uncertainty output quality of DINOZAUR and baseline models.
+Mean and standard deviation values are computed from three trained models.
+The best result is bold, second best is $\underline{\text{underlined}}$
 […]
-(×$(\times 10^{-2})$
+$(\times 10^{-2})$
 ShapeNet car
-GINOD $\underline{7.38}$ 18.58 0.36 16.64
-GINOL 10.38 $\underline{3.15}$ $\underline{0.27}$ 15.30
-TGINOD 9.18 54.13 0.41 23.17
-TGINOL 9.28 3.07 $\underline{0.27}$ $\underline{14.17}$
-DINOZAURD (Ours) 6.92 16.00 0.36 16.05
-DINOZAURL (Ours) 8.82 $\underline{3.15}$ $\underline{0.27}$ 15.30
-DINOZAURB (Ours) 7.46 4.35 0.20 12.73
+GINOD $7.94\pm 0.54$ $3.090\pm 0.044$ $0.288\pm 0.010$ $14.731\pm 0.625$
+GINOL $7.90\pm 0.29$ $3.106\pm 0.023$ $0.283\pm 0.007$ $14.709\pm 0.343$
+TGINOD $7.18\pm 0.17$ $2.996\pm 0.010$ $0.278\pm 0.005$ $13.704\pm 0.189$
+TGINOL $7.51\pm 0.73$ $3.071\pm 0.059$ $\underline{0.266\pm 0.018}$ $14.240\pm 0.811$
+DINOZAURD (Ours) $\mathbf{6.96\pm 0.04}$ $\underline{2.987\pm 0.003}$ $0.272\pm 0.002$ $\underline{13.479\pm 0.044}$
+DINOZAURL (Ours) $\underline{7.11\pm 0.05}$ $3.041\pm 0.008$ $\underline{0.266\pm 0.001}$ $13.749\pm 0.003$
+DINOZAURB (Ours) $7.49\pm 0.07$ $\mathbf{2.767\pm 0.011}$ $\mathbf{0.168\pm 0.001}$ $\mathbf{11.667\pm 0.136}$
 […]
@@
 In this section, we assess the UQ performance of the Bayesian version of our model, DINOZAURB, introduced in Section 3.3, on the more challenging non-uniform data.
-As baselines, we consider MC Dropout Gal and Ghahramani [2016] (NOD) and Laplace approximation Magnani et al.
-[2022] (NOL).
+As baselines, we consider MC Dropout [16] (NOD) and Laplace approximation [44] (NOL).
 MC Dropout is a widely used, model-agnostic UQ method, which we apply by enabling 10% dropout in the Fourier blocks’ MLPs to emulate stochasticity in the message passing.
-The Laplace approximation follows Magnani et al.
-[2022], modeling uncertainty via a fitted distribution over the parameters of the final MLP layer $\mathcal{Q}$.
+The Laplace approximation follows [44], modeling uncertainty via a fitted distribution over the parameters of the final MLP layer $\mathcal{Q}$.
 The discussion of hyperparameter choice, including $\mu_{\text{prior}},\sigma_{\text{prior}}$ required for setting $\ln\tau$ distributions is in the Appendix B.
-UQ results in Table 2 are obtained by estimating the mean and standard deviation from 100 posterior predictive distribution samples per test data instance.
-Metrics are computed using the Uncertainty Toolbox Chung et al.
-[2021]; full metric definitions are provided in the Appendix D.
-DINOZAURB matches or surpasses all baselines, achieving substantially lower negative log-likelihood (NLL) than dropout models and comparable results to Laplace approximation.
-It outperforms all baselines in terms of miscalibration area (MA) and interval score (IS).
+UQ metrics in Table 2 are calculated by drawing 100 posterior predictive distribution samples per test data instance.
+The reported mean and standard deviation are obtained from three independently trained models with random weight initialization.
+Metrics are computed using the Uncertainty Toolbox [61]; full metric definitions are provided in the Appendix D.
+Aleatoric noise was added at prediction time to all model outputs, using a fixed $\sigma^{2}=e^{-4}$ for dropout and Laplace models, and a learned variance term for Bayesian DINOZAUR that converged to $\sigma^{2}\approx e^{-4.8}$.
+DINOZAURB matches or surpasses all baselines, achieving lower negative log-likelihood (NLL) than dropout models or Laplace approximation.
+It also outperforms all baselines in terms of miscalibration area (MA) and interval score (IS).
 An important consideration is that the Laplace approximation requires a secondary computationally expensive step to obtain the Hessian of the loss function with respect to the network parameters after training.
@@
 To evaluate the scalability of DINOZAUR, we conduct experiments varying model width, depth, and the number of truncation modes using the ShapeNet car dataset.
-As shown in Figure 4(a), test error (measured by RL2 loss) decreases with increasing hyperparameters, saturating near the values used in our main experiments.
+As shown in Figure 4(a), test error (measured by RL2 loss) decreases with increasing architectural hyperparameters, saturating near the values used in our main experiments.
 Further improvements may be limited by the input data resolution of $64^{3}$.
@@
 Our solution is to include gradient features, though at the cost of increased computational time.
-Investigating other strategies, like introducing anisotropy intrinsically, using the modifications proposed in Liu-Schiaffini et al.
-[2024], or mixing gradients with other features before applying inverse FFT overhead, are promising directions for future work.
+Investigating other strategies, like introducing anisotropy intrinsically, using the modifications proposed in [62], or mixing gradients with other features before applying inverse FFT overhead, are promising directions for future work.
 We found that our Bayesian formulation is sensitive to hyperparameter settings of the prior distributions.
@@
 More broadly, exploring other parsimonious kernel parameterizations may offer both interpretability and improved generalization, especially in settings with limited data or strong physical priors.
-Finally, since our parameterization is independent of spatial dimensionality, it opens up opportunities for pretraining on lower-dimensional or simpler physical systems, with the potential to accelerate learning in more complex, higher-dimensional domains through transfer learning Hao et al.
-[2024].
+Finally, since our parameterization is independent of spatial dimensionality, it opens up opportunities for pretraining on lower-dimensional or simpler physical systems, with the potential to accelerate learning in more complex, higher-dimensional domains through transfer learning [63].
 We believe our work is a step forward to a deeper understanding of neural operators, advancing them toward becoming reliable tools for real-world scientific and engineering applications.
+Acknowledgments and Disclosure of Funding
+We want to thank Greg Bellchambers, Bachir Djermani, Axen Georget, Pavel Shmakov, and Phoenix Tse for insightful discussions and collaboration on the implementation.
+CRediT author statement: Albert Matveev: Conceptualization, Methodology, Software, Investigation, Writing - Original Draft, Visualization.
+Sanmitra Ghosh: Conceptualization, Methodology, Software, Investigation, Writing - Review and Editing, Visualization.
+Aamal Hussain: Software, Investigation, Writing - Original Draft.
+James-Michael Leahy: Conceptualization, Methodology, Software, Writing - Review and Editing.
+Michalis Michaelides: Conceptualization, Resources, Writing - Original Draft, Supervision, Project administration.
 Appendix A Benchmarks
@@
 Dataset $d$ Input type Output type $N_{\mathrm{train}},N_{\mathrm{test}}$
-Darcy flow 2 Regular [241;241][241;241][ 241 ; 241 ] Regular [241;241][241;241][ 241 ; 241 ] 1024, 1024
-NS 2D+T 2+1 Regular [64;64;10][64;64;10][ 64 ; 64 ; 10 ] Regular [64;64;10][64;64;10][ 64 ; 64 ; 10 ] 1000, 200
-ShapeNet car 3 Regular [64;64;64][64;64;64][ 64 ; 64 ; 64 ] Irregular [3,586][3{,}586][ 3 , 586 ] 561, 100
-[…]
+Darcy flow 2 Regular $[241;241]$ Regular $[241;241]$ 1024, 1024
+NS 2D+T 2+1 Regular $[64;64;10]$ Regular $[64;64;10]$ 1000, 200
+ShapeNet car 3 Regular $[64;64;64]$ Irregular [3,586][3{,}586] 561, 100
+Ahmed bodies 3 Regular $[64;64;64]$ Irregular $\sim[150{,}000]$ 500, 51
 Table 4: Data transformations during training
 Dataset Inputs Input scaler Padded input Target Target scaler
-Darcy flow Permeability Standard [271;271][271;271][ 271 ; 271 ] Pressure Standard
+Darcy flow Permeability Standard $[271;271]$ Pressure Standard
 […]
-ShapeNet car SDF None [72;72;72][72;72;72][ 72 ; 72 ; 72 ] Pressure Standard
-Ahmed bodies SDF, Inlet velocity None [72;72;72][72;72;72][ 72 ; 72 ; 72 ] Pressure Standard
+ShapeNet car SDF None $[72;72;72]$ Pressure Standard
+Ahmed bodies SDF, Inlet velocity None $[72;72;72]$ Pressure Standard
 To evaluate our model and compare it against the baselines, we selected two datasets with regularly sampled domains and two datasets with non-uniform outputs.
 Darcy flow
-This dataset contains steady-state 2D solutions of the Darcy flow equation on the unit box, modeling the flow of fluid through a porous medium Li et al.
-[2021].
-It includes samples with spatial resolution of $\times 241$ 241 points.
+This dataset contains steady-state 2D solutions of the Darcy flow equation on the unit box, modeling the flow of fluid through a porous medium [9].
+It includes samples with spatial resolution of $241\times 241$ points.
 The network received a scalar permeability field as input and was tasked to predict scalar pressure values.
-To ensure non-periodicity, we applied padding during training, increasing the resolution to $\times 271$ 271.
+To ensure non-periodicity, we applied padding during training, increasing the resolution to $271\times 271$.
 A standard scaler was applied to both input and output fields.
@@
 Navier-Stokes
-The Navier–Stokes dataset includes simulations of incompressible, viscous fluid flow with constant density and viscosity $\nu=10^{-5}$ Li et al.
-[2021].
-It features a periodic 2D domain of resolution $\times 64$ 64, with 20 time steps of the vorticity field evolution.
+The Navier–Stokes dataset includes simulations of incompressible, viscous fluid flow with constant density and viscosity $\nu=10^{-5}$ [9].
+It features a periodic 2D domain of resolution $64\times 64$, with 20 time steps of the vorticity field evolution.
 We trained models to predict the final 10 time steps from the initial 10, using 1000 training and 200 test samples.
 Inputs were composed as 3D tensors containing 10 time steps for every spatial point.
-Padding was introduced only to the time dimension, making the input to be of 64×64×1564$\times$ 64$\times$ 1564 × 64 × 15.
-We appended spatial coordinates uniformly sampled in $[0,2\pi)$ and a time coordinate with an integer index $t=[0,1,\dots 9]$, which resulted in 4 input channels that were mapped to one channel.
+Padding was introduced only to the time dimension, making the input to be of $64\times 64\times 15$.
+We appended spatial coordinates uniformly sampled in $0,2\pi)$ and a time coordinate with an integer index $t=[0,1,\dots 9]$, which resulted in 4 input channels that were mapped to one channel.
 ShapeNet car
-This benchmark contains 661 car meshes paired with Reynolds-Averaged Navier–Stokes (RANS) simulations Umetani and Bickel [2018].
+This benchmark contains 661 car meshes paired with Reynolds-Averaged Navier–Stokes (RANS) simulations [59].
 We focus on predicting the scalar pressure field at the mesh vertices, which are fixed in size at 3,586 points per mesh.
@@
 Ahmed bodies
-The Ahmed bodies dataset Ahmed et al.
-[1984] follows a similar setup to the ShapeNet car but introduces varying inlet velocities as an additional signal.
+The Ahmed bodies dataset [60] follows a similar setup to the ShapeNet car but introduces varying inlet velocities as an additional signal.
 This constant is appended to the SDF input, resulting in two input channels.
@@
 The grid resolution, padding, and target scaling match those used in the ShapeNet car.
-The GINO-decoder architecture Li et al.
-[2024] used for both ShapeNet and Ahmed datasets requires precomputed fixed-radius neighborhoods to map grid samples to mesh vertices.
-We use the radii from the original paper: 0.050.050.05 for the ShapeNet car and 0.0350.0350.035 for Ahmed bodies.
-For models trained on non-uniform data, we additionally apply Transformer-style positional encoding Vaswani et al. [2017] to the input features, grid locations, and mesh vertices.
+The GINO-decoder architecture [10] used for both ShapeNet and Ahmed datasets requires precomputed fixed-radius neighborhoods to map grid samples to mesh vertices.
+We use the radii from the original paper: $0.05$ for the ShapeNet car and $0.035$ for Ahmed bodies.
+For models trained on non-uniform data, we additionally apply Transformer-style positional encoding [64] to the input features, grid locations, and mesh vertices.
 Results in the main text are reported on the unscaled targets. Raw dataset specifications are provided in Table 3, and data transformations are detailed in Table 4.
@@
 GINOD 64 [24, 24, 24] 0.4 False
-TGINOD 64 [24, 24, 24] $10^{-4}$ False $\mathrm{D}=0.1$
+TGINOD 64 [24, 24, 24] $10^{-4}$ False $\mathrm{D}=0.1,\,\sigma^{2}=e^{-4}$
 DINOZAURD 64 [24, 24, 24] No True
 GINOL 64 [24, 24, 24] 0.4 False $\sigma_{H}^{2}=e^{-4},$
-TGINOL 64 [24, 24, 24] $10^{-4}$ False $C=10^{6},$
-DINOZAURL 64 [24, 24, 24] No True $\alpha=500$
+TGINOL 64 [24, 24, 24] $10^{-4}$ False $C=500,$
+DINOZAURL 64 [24, 24, 24] No True $\alpha=10^{6}$
 DINOZAURB 64 [24, 24, 24] No True $\mathbf{N}_{(\ln\tau)}(\ln 0.01,1),$
@@
 All models featured 4 FNO blocks with GELU activations.
-The channel width $d_{c}$ and the number of modes were selected to align with the original configurations of FNO Li et al.
-[2021] and GINO Li et al.
-[2024], with the exception of Navier-Stokes data.
+The channel width $d_{c}$ and the number of modes were selected to align with the original configurations of FNO [9] and GINO [10], with the exception of Navier-Stokes data.
 We increased modes and channels to reflect the complexity of this benchmark.
@@
 $H_{C,\alpha}=C\mathbb{E}\left[-\frac{\partial^{2}\log p(\mathcal{D}\mid\theta_{\mathcal{Q}})}{\partial\theta^{2}}\right]+\alpha I$
-where the expectation is approximated via Monte Carlo samples, and $C$ and $\alpha$ are hyperparameters selected on a validation set. In our experiments, we set $C=10^{6},\,\alpha=500$
+where the expectation is approximated via Monte Carlo samples, and $C$ and $\alpha$ are hyperparameters selected on a validation set. In our experiments, we set $C=500,\,\alpha=10^{6}$
 DINOZAURB requires us to set the prior distribution parameters for each $\ln\tau_{i}^{c}$ and specify the initial values for the approximate posterior $\ln\tau_{i}^{c}\mid\mathcal{D}$; we initialize this distribution with a full-rank covariance matrix within each NO block.
@@
 Optimization hyperparameters are detailed in Table 6.
-All models were trained using either AdamW Schedule-Free Defazio et al.
-[2024] or AdamW combined with a One-Cycle scheduler Smith and Topin [2019], weight decay set to $10^{-4}$.
+All models were trained using either AdamW Schedule-Free [57] or AdamW combined with a One-Cycle scheduler [56], weight decay set to $10^{-4}$.
 Both options allow for faster convergence with fewer epochs and reduce sensitivity to hyperparameter choices.
 The final setup was selected by comparing both options and reporting the results for the configuration that achieved the best performance.
-A limited learning rate sweep was conducted in the range $[10^{-4},10^{-2}]$, with early termination applied to underperforming runs.
+A limited learning rate sweep was conducted in the range $10^{-4},10^{-2}]$, with early termination applied to underperforming runs.
 For the One-Cycle scheduler, the learning rate in the table refers to the max LR parameter.
@@
 Figure 7: Lower diagonal full-rank learned covariance matrices of the posterior distributions.
-Each matrix represents a $d_{c}\times d_{c}$ ($\times 64$ 64) covariance matrix of the multi-dimensional distribution of diffusion times within each block.
+Each matrix represents a $d_{c}\times d_{c}$ ($64\times 64$) covariance matrix of the multi-dimensional distribution of diffusion times within each block.
 Finally, in Figure 7, we provide the full covariance matrices for each of the posteriors that were learned by the network.
@@
 For $i=1,\dots,M$, each block $\mathcal{N}_{\theta_{i},\tau_{i}}$ is of the form
-$\mathcal{N}_{\theta_{i},\tau_{i}}[v](x):=\sigma\left(W_{i}^{\mathrm{skip}}v(x)+b_{i}+W_{i}^{\mathrm{mix}}\begin{bmatrix}\mathcal{F}^{-1}\bigl{[}\exp(-4\pi^{2}\|k\|^{2}\tau_{i})\odot\mathcal{F}[v](k)\bigr{]}(x)\\ \mathcal{G}_{W_{i}^{\mathrm{grad}}}\Bigl{[}\mathcal{F}^{-1}\bigl{[}\exp(-4\pi^{2}\|k\|^{2}\tau_{i})\odot\mathcal{F}[v](k)\bigr{]}\Bigr{]}(x)\end{bmatrix}\right).$
+[…]
 Each hidden block defines a mapping $\mathcal{N}_{\theta_{i},\tau_{i}}:\mathcal{V}(\Omega,\mathbb{R}^{d_{c}})\rightarrow\mathcal{V}(\Omega,\mathbb{R}^{d_{c}})$.
@@
 $=\sigma\left(W_{1}^{\mathrm{skip}}v(x)+b_{1}+\fint_{\mathbb{T}^{d}}v(y)\mathrm{d}y\right),$
-making a full network $\mathcal{N}_{\theta,\infty}=\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\infty}\circ\mathcal{P}$ an averaging neural operator in the sense of Equation 2.3 in Lanthaler et al.
-[2023] on the torus $\mathbb{T}^{d}$.
-By Corollary 3.3 Lanthaler et al.
-[2023], $\mathcal{N}_{\theta,\infty}$ is a universal approximator on the torus.
-Here, we note that the lifting and projection operators defined in Lanthaler et al.
-[2023] have an explicit positional dependency, and thus, one should understand them as periodized MLPs.
-111 Note that Theorem 2.1 in Lanthaler et al.
-[2023] as it is formulated, applies to bounded Lipschitz domains.
-Thus, the proof of Corollary 3.3 Lanthaler et al.
-[2023] is not immediate.
+making a full network $\mathcal{N}_{\theta,\infty}=\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\infty}\circ\mathcal{P}$ an averaging neural operator in the sense of Equation 2.3 in [66] on the torus $\mathbb{T}^{d}$.
+By Corollary 3.3 [66], $\mathcal{N}_{\theta,\infty}$ is a universal approximator on the torus.
+Here, we note that the lifting and projection operators defined in [66] have an explicit positional dependency, and thus, one should understand them as periodized MLPs.
+[^1] Note that Theorem 2.1 in [66] as it is formulated, applies to bounded Lipschitz domains.
+Thus, the proof of Corollary 3.3 [66] is not immediate.
 One must either follow the proof of Theorem 2.1 to see that it applies to the torus, or first formulate it on the domain $(0,2\pi)^{d}$, invoke the Theorem 2.1, and then argue that it holds when one restricts to periodic functions on $(0,2\pi)^{d}$.
@@
 By Lemma 41 by Kovachki et al.
-[2021] and Lemma A1 in Lanthaler et al.
-[2023], there exists a continuous, linear extension operator $\mathcal{E}:C^{s}(\Omega;\mathbb{R}^{d_{a}})\rightarrow C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})$ such that $\mathcal{E}[a]|_{\Omega}=a$ and $\mathcal{E}[a]$ is periodic on $[0,2\pi]^{d}$ for all $a\in C^{s}(\Omega;\mathbb{R}^{d_{a}})$.
-We note that the conditions of the Theorem 9 in Kovachki et al.
-[2021] are satisfied.
+[67] and Lemma A1 in [66], there exists a continuous, linear extension operator $\mathcal{E}:C^{s}(\Omega;\mathbb{R}^{d_{a}})\rightarrow C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})$ such that $\mathcal{E}[a]|_{\Omega}=a$ and $\mathcal{E}[a]$ is periodic on $0,2\pi]^{d}$ for all $a\in C^{s}(\Omega;\mathbb{R}^{d_{a}})$.
+We note that the conditions of the Theorem 9 in [67] are satisfied.
 By following the proof of this theorem, we apply the extension operator $\mathcal{E}$ and use the universal approximation on the torus, establishing universal approximation on $\Omega$.